### v2

# Lean Formalization of Generalization Error Bound by Rademacher Complexity

Abstract
We formalize the generalization error bound using Rademacher complexity in the Lean 4 theorem prover. Generalization error quantifies the gap between a learning machine’s performance on given training data versus unseen test data, and Rademacher complexity serves as an estimate of this error based on the complexity of learning machines, or hypothesis class. Unlike traditional methods such as PAC learning and VC dimension, Rademacher complexity is applicable across diverse machine learning scenarios including deep learning and kernel methods. We formalize key concepts and theorems, including the empirical and population Rademacher complexities, and establish generalization error bounds through formal proofs of McDiarmid’s inequality, Hoeffding’s lemma, and symmetrization arguments. We make our code available on GitHub. [^1] https://github.com/auto-res/lean-rademacher
1 Introduction
Generalization is a central concept in machine learning that describes how well a learning machine, constructed based on training data, can make predictions on unseen test data. In practice, minimizing the training error is desirable, but this alone does not necessarily guarantee a better performance on test error. Overfitting occurs when a machine excessively fits the training data, leading to poor predictive performance on test data. The deviation between the training and test errors is called the generalization error, or the generalization gap. To quantitatively evaluate this issue and ensure the reliability of learning results, theoretical analysis of generalization error is essential.
In this study, we explain the generalization error bound using Rademacher complexity [4] and presents its formalization in Lean 4 [5]. The Rademacher complexity measures the size of a set in a sense, and is a de-facto standard method to derive upper bounds on generalization error in the modern machine learning problems. Formally, it quantifies how well a hypothesis class can fit random noise, represented by Rademacher variables. Rademacher complexity is broadly applicable to both classification and regression problems, as it can be defined for arbitrary bounded loss functions. Unlike PAC learning [13] and VC dimension [14], which provide worst-case uniform bounds over all datasets, Rademacher complexity gives a finer, data-dependent estimate of model complexity. This makes it a powerful tool for analyzing modern machine learning models, including deep learning and kernel methods [8, 10, 2].
2 Statistical Machine Learning and Generalization Error
We formulate the concepts of statistical machine learning and generalization error. We refer to Mohri et al. (2018) [8] as a standard textbook on statistical machine learning theory, and Wainwright (2019) [7] for more mathematical details on the Rademacher complexity and concentration inequalities used in generalization error analysis.
Figure 1: Generalization error of a learning algorithm $A:\mathcal{X}^{n}\to\mathcal{F}$ refers to the gap between the original data source (concept $c\in\mathcal{C}$) and the estimated data generator (hypothesis $f\in\mathcal{F}$), which is an output of the learning algorithm $A$ given a dataset $\bm{x}\in\mathcal{X}^{n}$ generated according to concept $c$. Following the convention of graphical models, filled-in circles are observable, while unfilled circles are unobservable.
2.1 Conventions in Measure-Theoretic Probability
A probability space $(\Omega,\mathcal{M},\mu)$ is a triple composed of measurable space $\Omega$, $\sigma$-algebra $\mathcal{M}$, and finite positive measure $\mu$ satisfying $\mu(\Omega)=1$.
A random variable $X:\Omega\to\mathcal{X}$ is a measurable map from $\Omega$ to a measurable space $\mathcal{X}$. A realization $x\in\mathcal{X}$ of a random variable $X$ is an image $X(\omega)$ of $X$ at a certain element $\omega\in\Omega$. Following the convention, we use the uppercase for random variables (e.g. $X,Y,\ldots$), while the lowercase for their realizations (e.g. $x,y,\ldots$). The distribution (or law) $P$ of a random variable $X$ is the push-forward measure $X_{\sharp}\mu$ of $\mu$ by $X$. We write $X\sim P$, and say $X$ is a random sample in $\mathcal{X}$ drawn according to probability distribution $P$, when $X$ is a random variable with distribution $P$.
We say a sequence $\bm{X}:=\{X_{k}\}_{k=1}^{n}$ of random variables (or a random vector for short) is independent and identically distributed (i.i.d.), denoted $\bm{X}\sim^{iid}P$, when each $X_{k}$ has the same distribution $P$, and all are mutually independent, so as the distribution $P_{\bm{X}}$ of $\bm{X}$ is given by the product $P_{\bm{X}}(\bm{X})=P^{n}(\bm{X}):=\prod_{k=1}^{n}P(X_{k})$.
The expectation of a measurable function $f:\mathcal{X}\to\mathbb{R}$, denoted $\mathbb{E}_{X}[f(X)]$ or $P[f(X)]$, is the integration of $f$ with respect to the measure $P$, that is, $\mathbb{E}_{X}[f(X)]:=\int_{\mathcal{X}}f(x)\mathrm{d}P(x)=\int_{\Omega}f\circ X(\omega)\mathrm{d}\mu(\omega)$. By $L^{1}(P)$, we write the Banach space of all $P$-integrable functions $f:\mathcal{X}\to\mathbb{R}$.
Those concepts are formalized in namespace MeasureTheory in Mathlib. Especially, in our source code, the probability space $(\Omega,\mathcal{M},\mu)$ and a random vector $\bm{X}=\{X_{k}\}_{k=1}^{n}:\Omega\to\mathcal{X}^{n}$ are typically formalized as follows.
universe u v
variable : Type u} [MeasurableSpace Ω]
variable : Measure Ω} [IsProbabilityMeasure μ]
variable {甲 : Type v} [MeasurableSpace 甲]
variable {n : ℕ} {X : Fin n Ω 甲} (hX : k : Fin n, Measurable (X k))
2.2 Formulation of Statistical Machine Learning
The sample space $\mathcal{X}$ is a measurable set of datasets. Following convention, an observation (an example, or a datum) refers to a single element $x\in\mathcal{X}$, while a dataset (a sample, or a data) refers to a single sequence of observations $\bm{x}=\{x_{k}\}_{k=1}^{n}\subset\mathcal{X}^{n}$. For example, in the image recognition problem, a single dataset $\bm{x}\in\mathcal{X}^{n}$ is an $n$-fold pairs $\{(\mathtt{image}_{k},\mathtt{label}_{k})\}_{k=1}^{n}$ of images and labels. In statistical machine learning, a single dataset $\bm{x}=\{x_{k}\}_{k=1}^{n}$ is formulated as a realization of a random vector $\bm{X}=\{X_{k}\}_{k=1}^{n}:\Omega\to\mathcal{X}^{n}$. In the following, we omit emphasizing the dependency in sample size $n$, and simply write $\bm{x},\bm{X},\mathcal{X}^{n}$ as $x,X,\mathcal{X}$ if not necessary.
The concept class $\mathcal{C}$ is a collection of data sources (called concepts) that describe how datasets are obtained. In this study, we assume $\mathcal{C}$ to be a family of random vectors $\bm{X}:\Omega\to\mathcal{X}^{n}$, or equivalently, probability distributions $P$ on $\mathcal{X}^{n}$. In the example of the image recognition problem, a single concept is a random vector $\bm{X}:\Omega\to\mathcal{X}^{n}$, or its law $P_{\bm{X}}(\{\mathtt{image}_{k},\mathtt{label}_{k}\}_{k=1}^{n})$. In statistical machine learning, the concept itself is supposed to be unobservable, and only the dataset to be observable. In other words, only an image $\bm{x}=\bm{X}(\omega)$ is given, but the map $\bm{X}:\Omega\to\mathcal{X}^{n}$ itself is not given.
The hypothesis class $\mathcal{F}$ is another collection of data generators (called hypotheses, or learning machines). Like concepts, hypotheses describe the data generation process. However, unlike concepts, hypotheses have parameters, say $\theta\in\Theta$, that we can freely manipulate. For example, in deep learning, the hypothesis class is a set of deep neural networks (DNNs).
A learning algorithm $A$ is a measurable map $\mathcal{X}^{n}\to\mathcal{F}$ that describes how to associate datasets with hypotheses. Regarding the data generation process by concepts as a forward process, learning algorithm $A$ corresponds to a backward process. In the terms of statistical estimation theory, $A$ is an estimator, and the learned machine $f$ is an image $f=A(\bm{x})$ of a given dataset $\bm{x}\in\mathcal{X}^{n}$. For example, in deep learning, $A$ is the process of empirical risk minimization by using stochastic gradient descent on the parameter space of DNNs.
As illustrated in Figure 1, the generalization error (explained in the next subsection) estimates the discrepancy between the learned hypothesis $f=A(\bm{x})$ and the original concept $c$. A learning algorithm is considered better if its generalization error is smaller.
In our source code, especially, a hypotheses class =$\mathcal{F}=\{f_{i}:\mathcal{X}\to\mathbb{R}\mid i\in\iota\}$ is typically formalized as follows.
universe w
variable : Type w} {f: ι ℝ} (hf : i, Measurable (f i))
2.3 Formulation of Generalization Error Analysis
The training data(set) refers to a random vector $\bm{X}:\Omega\to\mathcal{X}^{n}$, and a test data(set) refers to another random variable $X^{\prime}:\Omega\to\mathcal{X}$ that is statistically independent from training data $\bm{X}$. In this study, we assume (1) that the training dataset is i.i.d., so $\bm{X}\sim^{iid}P$, and (2) that both training and test datasets have the same distribution $P$, so $X^{\prime}\sim P$. We note that $P$ itself is unknown, although we are supposed to know that $\bm{X}$ and $X^{\prime}$ have the same distribution. This may sound technical, but natural when i.i.d. sampling is easy, thus often assumed in the basic formulations.
A (pointwise) loss function :×$\ell:\mathcal{F}\times\mathcal{X}\to\mathbb{R}_{\geq 0}$ is a measurable functional that associates hypotheses with positive numbers. For example, the squared error loss $\ell(f,(x,y)):=|f(x)-y|^{2}$ is typical in supervised learning.
The training error (a.k.a. empirical risk) $L(f\mid\bm{X})$ of a hypothesis $f$ over a training dataset $\bm{X}$ is the sample average of the pointwise loss function: $L(f\mid\bm{X}):=\frac{1}{n}\sum_{k=1}^{n}\ell(f,X_{k})$. We note that in the real-world application, we can only compute its realization, say $L(f\mid\bm{x})$.
The test error (a.k.a. population risk) $L(f)$ of a hypothesis $f$ (over a test dataset $X$) is the expectation of the pointwise loss function: $L(f):=\mathbb{E}_{X}[\ell(f,X)]$. Following the convention, the dependency on $X$ is omitted for simplicity. By the assumptions that both training and test datasets are i.i.d. samples, namely $(\bm{X},X)\sim^{iid}P$, the expectation of the training error over the i.i.d. draw of training dataset is identical to the test error: $\mathbb{E}_{\bm{X}}[L(f\mid\bm{X})]=L(f)$.
The test error $L(f)$ is understood as measuring the generalization performance of a hypothesis on unseen data, because the distribution $P$ of test data is unknown. Indeed, it is the (first) definition of the generalization error as explained soon below.
The generalization error refers to three related quantities: (1) population risk (or test error) $L(f)$ itself, (2) generalization gap (the gap between test and training errors) $\Delta(f):=L(f)-L(f\mid\bm{X})$, and (3) excess risk (the population risk relative to its infimum) $L(f)-\inf_{f\in\mathcal{F}}L(f)$. In all three definitions, the interest lies in the (either absolute or relative) value of population risk $L(f)$, and either one will be obtained depending on the estimation technique employed.
In this study, the main theorem (Theorem 1) presents an upper bound on generalization gap $\Delta(f)$ by using the Rademacher complexity, which estimates the second meaning of generalization error. We note that, as clarified in the remark (Remark 2), an upper bound of the gap $L(f)-L(f\mid\bm{X})\leq B$ can be trivially turned into the upper bound of the risk $L(f)\leq L(f\mid\bm{X})+B$, which estimates the first meaning of generalization error.
3 Rademacher Complexity
We define the empirical and population Rademacher complexities of a hypothesis class $\mathcal{F}$.
Definition 1 (Rademacher Variable).
A uniform random variable $\sigma$ taking values in {±$\{\pm 1\}$ is called a Rademacher variable, and an i.i.d. sequence of Rademacher variables $\bm{\sigma}:=\{\sigma_{k}\}_{k=1}^{n}$ (i.e. uniform random vector taking values in {±$\{\pm 1\}^{n}$) is called a Rademacher vector.
Definition 2 (Rademacher Complexity).
Let $\mathcal{F}\subset L^{1}(P)$ be a separable subspace of real-valued integrable functions on $\mathcal{X}$. Let $\bm{X}=\{X_{k}\}_{k=1}^{n}$ be an i.i.d. random vector drawn from distribution $P$. The Empirical Rademacher complexity is defined as
$\mathcal{R}(\mathcal{F}\mid\bm{X}):=\mathbb{E}_{\bm{\sigma}}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}\sigma_{k}f(X_{k})\Bigg{|}\right]=\frac{1}{2^{n}}\sum_{\bm{\sigma}\in\{\pm 1\}^{n}}\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}\sigma_{k}f(X_{k})\Bigg{|},$ (1)
and the (population) Rademacher complexity is defined as its expectation
$\mathcal{R}_{n}(\mathcal{F}):=\mathbb{E}_{\bm{X}}[\mathcal{R}(\mathcal{F}\mid\bm{X})]=\int_{\Omega^{n}}\left[\frac{1}{2^{n}}\sum_{\bm{\sigma}\in\{\pm 1\}^{n}}\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}\sigma_{k}f(X_{k}\circ\omega)\Bigg{|}\right]\mathrm{d}\mu^{n}(\omega).$ (2)
def Signs (n : ℕ) : Type := Fin n ({-1, 1} : Finset ℤ) Rademacher vector
noncomputable
def empiricalRademacherComplexity
(n : ℕ) (f : ι ℝ) (x : Fin n 甲) : :=
(Fintype.card (Signs n) : ℝ)⁻¹ *
σ : Signs n, i, |(n : ℝ)⁻¹ * k : Fin n, k : ℝ) * f i (x k)|
noncomputable
def rademacherComplexity
(n : ℕ) (f : ι ℝ) : Measure Ω) (X : Ω 甲) : :=
μⁿ[fun ω : Fin n Ω empiricalRademacherComplexity n f (X ω)]
Here $\mathtt{\mu^{n}}$ is a local notation for the product measure $\mu^{n}$ defined as below.
local notation "μⁿ" => Measure.pi (fun _ μ)
4 Generalization Error Bounds by Rademacher Complexity
Here, we explain the main theorem and its formalization. We refer to Theorem 4.10 from Wainwright (2019) [7] and Theorem 3.3 from Mohri et al. (2018) [8]. To be precise, Mohri et al.’s statement is a corollary of Wainwright’s statement tailored for practical purpose. So, we explain Wainwright’s version as the main theorem, and Mohri et al.’s version in the remark.
Theorem 1 (Generalization Error Bound by Rademacher Complexity).
Suppose that the hypothesis class $\mathcal{F}$ (includes the loss function and) is $b$-uniformly bounded, namely there exists a scalar $b\geq 0$ such that $\sup_{f\in\mathcal{F}}\|f\|_{L^{\infty}(\mathcal{X})}\leq b$. For any positive integer $n\geq 1$ and scalar $\varepsilon\geq 0$, the following holds with probability at least $1-\exp\left(-\frac{n\varepsilon^{2}}{2b^{2}}\right)$:
$\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}f(X_{k})-\mathbb{E}_{X}[f(X)]\Bigg{|}\leq 2\mathcal{R}_{n}(\mathcal{F})+\varepsilon$ (3)
Here, the phrase “An event $E(\varepsilon)$ with parameter $\varepsilon$ holds with probability at least $1-\beta(\varepsilon)$” is a predicate of probability theory, meaning that the following inequality (a.k.a. concentration of probability measure, or tail probability bound, with rate function $\beta$) for complement event $E^{c}(\varepsilon)$ holds:
$P(E^{c}(\varepsilon))\leq\beta(\varepsilon),$ (4)
which implies $P(E(\varepsilon))\geq 1-\beta(\varepsilon)$. Hence, the statement is formalized as below.
theorem main
[MeasurableSpace 甲][Nonempty 甲][Nonempty ι][Countable ι][IsProbabilityMeasure μ]
(f : ι ℝ) (hf : i, Measurable (f i))
(X : Ω 甲) (hX : Measurable X)
{b : ℝ} (hb : 0 b) (hf : i x, |f i x| b)
{t : ℝ} (ht : 0 t) (ht : t * b ^ 2 1 / 2)
: ℝ} (hε : 0 ε) :
(μⁿ (fun ω 2 rademacherComplexity n f μ X + ε
uniformDeviation n f μ X (X ω))).toReal
(- ε ^ 2 * t * n).exp := by
Remark 1.
In the main theorem, we assume that a hypothesis includes a loss function. Namely, in the example of image recognition, a hypothesis is not a predictor $g:\mathcal{X}_{\mathtt{image}}\times\Theta\to\mathcal{X}_{\mathtt{label}}$ alone, but a composite $f:\mathcal{X}\times\Theta\to\mathbb{R}_{\geq 0},f((\mathtt{image},\mathtt{label}),\theta):=\ell(g(\mathtt{image},\theta),\mathtt{label})$ of $g$ followed by a loss function $\ell$ such as a cross-entropy or a squared error loss.
Remark 2.
In practice, the quantity of primary interest is the population risk $\mathbb{E}_{X}[\widehat{f}(X)]$ of the hypothesis $\widehat{f}=A(\bm{X})$ obtained by learning algorithm $A$. Because we are only given the training dataset $\bm{X}(\omega)\in\mathcal{X}$ (as a realization), and we do not know the data distribution $P_{\bm{X}}$ itself, this expectation is intractable. Nonetheless, as a consequence of the main theorem, the population risk can be estimated in a tractable manner as follows
$\mathbb{E}_{X}[\widehat{f}(X)]\leq\frac{1}{n}\sum_{k=1}^{n}\widehat{f}(X_{k})+2\mathcal{R}_{n}(\operatorname{\mathrm{Im}}A)+\sqrt{\frac{2b^{2}\log 1/\delta}{n}}$ (5)
with probability at least $1-\delta$ over the draw of an i.i.d. sample $\bm{X}$. This is Mohri et al’s version of the main theorem (Theorem 3.3 in [8]). We note that we further need to compute the Rademacher complexity separately depending on the specific problem.
Proof.
The proof is two-fold: Use McDiarmid’s (bounded difference) inequality, and symmetrization argument. Put the supremum $\Delta(\mathcal{F}\mid\bm{X})$ of the absolute deviation and its expectation $\Delta(\mathcal{F})$ as follows:
$\Delta(\mathcal{F}\mid\bm{X}):=\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}f(X_{k})-\mathbb{E}_{X}[f(X)]\Bigg{|},\quad\mbox{and}\quad\Delta(\mathcal{F}):=\mathbb{E}_{\bm{X}}[\Delta(\mathcal{F}\mid\bm{X})].$ (6)
We call $\Delta(\mathcal{F}\mid\bm{X})$ the uniform deviation for short, and formalize it as below.
def uniformDeviation (n : ℕ) (f : ι ℝ) : Measure Ω) (X : Ω 甲) : (Fin n 甲)
:= fun y i, |(n : ℝ)⁻¹ * k : Fin n, f i (y k) - μ [fun ω’ f i (X ω’) ]|
By McDiarmid’s inequality, the deviation of the uniform deviation from its mean is upper bounded by $\varepsilon$ as
$\Delta(\mathcal{F}\mid\bm{X})-\Delta(\mathcal{F})<\varepsilon$ (7)
with probability at least $1-\exp\left(-\frac{n\varepsilon^{2}}{2b^{2}}\right)$. In other words, the following inequality holds:
$\mu\left\{\omega\mid\Delta(\mathcal{F}\mid\bm{X})(\omega)-\Delta(\mathcal{F})\geq\varepsilon\right\}\leq\exp\left(-\frac{n\varepsilon^{2}}{2b^{2}}\right).$ (8)
theorem uniformDeviation_mcdiarmid
[MeasurableSpace 甲][Nonempty 甲][Nonempty ι][Countable ι][IsProbabilityMeasure μ]
{X : Ω 甲} (hX : Measurable X) (hf : i, Measurable (f i))
{b : ℝ} (hb : 0 b) (hf’: i x, |f i x| b)
{t : ℝ} (ht : 0 t) (ht : t * b ^ 2 1 / 2)
: ℝ} (hε : 0 ε) :
(μⁿ (fun ω : Fin n Ω uniformDeviation n f μ X (X ω) -
μⁿ[fun ω : Fin n Ω uniformDeviation n f μ X (X ω)] ε)).toReal
(- ε ^ 2 * t * n).exp := by
Based on the symmetrization argument, estimate $\Delta(\mathcal{F})$ by the Rademacher complexity as below. Take another i.i.d. sequence $\bm{Y}:=\{Y_{k}\}_{k=1}^{n}\sim^{iid}P$ independent of $\bm{X}$. Then,
$\Delta(\mathcal{F})$ $=\mathbb{E}_{\bm{X}}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}\left\{f(X_{k})-\mathbb{E}_{Y_{k}}[f(Y_{k})]\right\}\Bigg{|}\right]$ (9)
$=\mathbb{E}_{\bm{X}}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\mathbb{E}_{\bm{Y}}\left[\frac{1}{n}\sum_{k=1}^{n}\left\{f(X_{k})-f(Y_{k})\right\}\right]\Bigg{|}\right]$ (10)
$\leq\mathbb{E}_{\bm{X},\bm{Y}}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}\left\{f(X_{k})-f(Y_{k})\right\}\Bigg{|}\right]$ (11)
$=\mathbb{E}_{\bm{X},\bm{Y},\bm{\sigma}}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}\sigma_{i}\left\{f(X_{k})-f(Y_{k})\right\}\Bigg{|}\right]$ (12)
$\leq 2\mathbb{E}_{\bm{X},\bm{\sigma}}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}\sigma_{i}f(X_{k})\Bigg{|}\right]=2\mathcal{R}_{n}(\mathcal{F}).$ (13)
theorem le_two_smul_rademacher
[Nonempty ι][Countable ι][IsProbabilityMeasure μ]
(X : Ω 甲) (hf : i, Measurable (f i X))
{b : ℝ} (hb : 0 b) (hf’: i x, |f i x| b) :
μⁿ[fun ω : Fin n Ω uniformDeviation n f μ X (X ω)]
2 rademacherComplexity n f μ X := by
Finally, the combination of the estimates (7) and (13) yields the assertion. ∎
5 McDiarmid’s Inequality
McDiarmid’s inequality, a.k.a. the bounded difference inequality, estimates the concentration bound of a function $f:\mathcal{X}^{n}\to\mathbb{R}$ around its mean $\mathbb{E}[f(\bm{X})]$ under the assumption that $f$ satisfies the bounded difference property. We note that the case when $f$ is a sum $f(\bm{X})=\sum_{k=1}^{n}X_{k}$ reproduces Hoeffding’s inequality. We refer to Corollary 2.21 from [7] for more details.
There are two ways to prove this inequality: directly by using Hoeffding’s lemma (explained later), or indirectly as a corollary of the Azuma-Hoeffding inequality (see e.g. Corollary 2.20 in [7]). In this study, we employed the former direct way. We remark that the Azuma-Hoeffding inequality has never been formalized, which itself is an important future work as well [^2] It was implemented in Mathlib.Probability.Moments.SubGaussian at the same time of the first submission of this draft..
Definition 3 (Bounded Difference Property).
Given an $n$-tuple $\bm{x}\in\mathcal{X}^{n}$ and an element $x^{\prime}\in\mathcal{X}$, let $\bm{x}(k,x^{\prime})$ denote a new $n$-tuple obtained by replacing the $k$-th component $x_{k}$ of $\bm{x}$ with $x^{\prime}$. A function $f:\mathcal{X}^{n}\to\mathbb{R}$ satisfies the bounded difference property if there exists a sequence $\{c_{k}\}_{k=1}^{n}$ of positive numbers such that for all $k\in[n],\bm{x}\in\mathcal{X}^{n},x^{\prime}\in\mathcal{X}$:
$|f(\bm{x}(k,x^{\prime}))-f(\bm{x})|\leq c_{k}.$ (14)
Theorem 2 (One-sided McDiarmid’s Inequality, or Bounded Differences -).
Suppose that a measurable function $f:\mathcal{X}^{n}\to\mathbb{R}$ satisfies the bounded difference property with bounds $\{c_{k}\}_{k=1}^{n}$, and suppose that a real number $t\in\mathbb{R}$ satisfies $t\leq 1/\sum_{k=1}^{n}c_{k}^{2}$. Let $\bm{X}=\{X_{k}\}_{k=1}^{n}\sim P$ be an i.i.d. sequence. Then, for any $\varepsilon\geq 0$, we have:
$\mu\left\{\omega\mid f(\bm{X})(\omega)-\mathbb{E}_{\bm{X}}[f(\bm{X})]\geq\varepsilon\right\}\leq\exp\left(-2\varepsilon^{2}t\right)$ (15)
theorem mcdiarmid_inequality_pos
(X : ι Ω 甲) (hX : i, Measurable (X i)) (hX : iIndepFun X μ)
(f : 甲) ℝ) (hf : Measurable f)(c : ι ℝ)
(hf : (i : ι) (x : ι 甲) (x : 甲),
|f x - f (Function.update x i x’)| c i) bounded difference property
: ℝ) (hε : ε > 0)(t : ℝ) (ht : t * i, (c i) ^ 2 1) :
(fun ω : Ω (f (Function.swap X)) ω
- μ[f (Function.swap X)] ε)).toReal
(-2 * ε ^ 2 * t).exp := by
6 Hoeffding’s Lemma
Hoeffding’s lemma states that an almost surely bounded random variable $X$ is sub-Gaussian. It is used to show Hoeffding’s inequality and its generalization McDiarmid’s inequality. We refer to Lemma D.1 in [8] for more details. In the proof, we use exponential tilting, which has already been implemented in Mathlib.Probability.Moments.Tilted.
Theorem 3 (Hoeffding’s Lemma).
For a real random variable $X$ with $\mathbb{E}[X]=0$ and $X\in[a,b]$ almost surely, the inequality
$\mathbb{E}_{X}\left[\exp tX\right]\leq\exp\left(\frac{t^{2}(b-a)^{2}}{8}\right)$ (16)
holds almost surely for all $t\in\mathbb{R}$.
theorem hoeffding [IsProbabilityMeasure μ]
(t a b : ℝ) {X : Ω ℝ} (hX : AEMeasurable X μ)
(h : ∀ᵐ ω ∂μ, X ω Set.Icc a b) (h0 : μ[X] = 0) :
mgf X μ t exp (t^2 * (b - a)^2 / 8) := by
Here, mgf is the moment generating function $\mathbb{E}_{X}[\exp tX]$ of a real random variable $X$ defined in Mathlib. Probability.Moments.Basic as below:
def mgf (X : Ω ℝ) : Measure Ω) (t : ℝ) : := μ[fun ω => exp (t * X ω)]
7 Related Work
7.1 Formalization of Machine Learning Theory
Bagnall and Stewart (2019) [3] have also formalized a generalization error bound in Coq via Hoeffding’s inequality, which is limited to finite hypothesis class, namely $|H|<\infty$. On the other hand, we have formalized a more versatile bound in Lean 4 via Rademacher complexity, which covers infinite hypothesis class.
Bagnall and Stewart (2019) [3] Coq Generalization error bounds for finite hypothesis class
Tassarotti et al. (2021) [9] Lean PAC learnability of decision stumps
Vajjha et al. (2021) [11] Coq Convergence of reinforcement learning algorithms
Vajjha et al. (2022) [12] Coq Stochastic approximation theorem
(ours) Lean Generalization error bound by Rademacher complexity
7.2 Formalization of Concentration Inequalities
As listed below, a few fundamental concentration inequalities were formalized in the 2020s in several languages. Compared to the advanced progress in Isabelle/HOL, there remains significant room for development in Lean and Coq.
Markov/Chebyshev Lean (Mathlib), Coq (MathComp-Analysis[1], IBM/FormalML[11]),
Isabelle/HOL (HOL-Probability)
Azuma-Hoeffding Lean (Mathlib)
McDiarmid (ours), Isabelle/HOL (AFP) [6]
8 Future Direction
In this study, we have formalized the generalization error bound by Rademacher complexity. The main theorem is a starting point for generalization error analysis for modern machine learning problems. Computing Rademacher complexity for specific problem settings, such as deep learning, kernel machines, and ridge/Lasso regression, is an important future work.
Acknowledgments
We would like to thank the mathlib community members, particularly Rémy Degenne and Yaël Dillies, for their valuable comments regarding formalization of Hoeffding’s lemma. We would like to thank Reynald Affeldt, Alessandro Bruni, Tetsuya Sato, Hiroshi Unno, and Yoshihiro Mizoguchi for having productive discussion on our formalization project. This work was supported by JSPS KAKENHI 20K03657, JST PRESTO JPMJPR2125, and JST Moonshot R&D Program JPMJMS2236.