### v3

# Lean Formalization of Generalization Error Bound by Rademacher Complexity

Abstract
We formalize the generalization error bound using the Rademacher complexity for the Lean 4 theorem prover based on the probability theory in the Mathlib 4 library. Generalization error quantifies the gap between a learning machine’s performance on given training data versus unseen test data, and the Rademacher complexity is a powerful tool to upper-bound the generalization error of a variety of modern learning problems. Previous studies have only formalized extremely simple cases such as bounds by parameter counts and analyses for very simple models (decision stumps). Formalizing the Rademacher complexity bound, also known as the uniform law of large numbers, requires substantial development and is achieved for the first time in this study. In the course of development, we formalize the Rademacher complexity and its unique arguments such as symmetrization, and clarify the topological assumptions on hypothesis classes under which the bound holds. As an application, we also present the formalization of generalization error bound for $L^{2}$-regularization models.
The code is available at https://github.com/auto-res/lean-rademacher.
1 Introduction
Generalization is a central concept in machine learning that describes how well a learning machine can make predictions on not only training data but also on unseen test data. In practice, minimizing the training error is desirable, but this alone does not necessarily guarantee a better performance on test error. When a machine excessively fits the training data, overfitting occurs, leading to poor predictive performance on test data. The deviation between the training and test errors is called the generalization error. To quantitatively estimate the generalization error and ensure the reliability of learning results, statistical learning theory studies the theoretical estimates of generalization error, or the generalization error bounds.
In this study, we explain the generalization error bound using Rademacher complexity [4] and presents its formalization in Lean 4 [7] based on probability theory formalized in Lean 4’s mathematical library, Mathlib 4 [15]. The Rademacher complexity measures the complexity of learning machines, and is a de-facto standard tool to upper-bound generalization errors in the modern machine learning problems. For example, the classical Vapnik-Chervonenkis (VC) dimension [19] for the PAC learning [18] setting can only provide data-independent worst-case uniform bounds for 0–1 classification problems, the Rademacher complexity can provide a sharper, data-dependent bounds for a variety of learning problems, including kernel methods and deep learning [12, 14, 2].
With the rapid growth of machine learning, the amount of papers associating theoretical generalization analysis has also been increasing. However, proofs of generalization bounds are typically long and involve complicated dependencies among assumptions, so even experts can verify only a limited number of such proofs by hand. We therefore expect increasing automation of generalization analysis via formal proofs. This work formalizes the “fundamental theorem” that, in a future where such automated generalization analysis is standard, is expected to be most frequently used in practice.
Indeed, the main theorem we present is a fundamental result also known as the Uniform Law of Large Numbers (ULLN), with broad applications across probability theory and mathematical statistics, beyond learning theory.
Modern machine learning problem settings range widely—from classical linear regression and binary classification to kernel methods, generative tasks with diffusion models, and in-context learning with large language models—so diverse formalizations tailored to each setting are required. Rademacher complexity is a crucial starting point for these efforts and merits a full-scale formalization.
Prior work has formalized generalization bounds, but only in relatively simplified settings. For example, Bagnall and Stewart (2019) [3] formalize a generalization bound for neural networks, but restrict parameters to a finite set, which is insufficient for formalizing today’s diverse generalization analyses. Our contribution is, to our knowledge, the first systematic formalization of generalization analysis via Rademacher complexity. As an application of the general theorem, we also formalize the generalization error bound of $L^{2}$-regularization models.
In formalizing statistical learning theory, we adopted the Lean theorem prover [7]. This is mainly because authors were relatively familiar with Lean, and Mathlib library’s [15] measure-theoretic probability was well developed. For details on Mathlib’s probability theory, we refer to the explanatory blog post by the developer himself [8].
2 Preliminaries
To formalize the “generalization error,” we first overview measure-theoretic probability, mathematical statistics, and statistical learning theory. For historical reasons, statistical learning theory—and its foundations in mathematical statistics and measure-theoretic probability—contain many domain-specific terms that often cause confusion. For example, terms such as “data,” “model,” “random,” “error,” “hypothesis,” “concept,” and “risk” are defined differently than in neighboring fields such as numerical analysis and physics. Fortunately, measure-theoretic probability is well developed in the Mathlib library. Thus, to formalize generalization error, here we first translate all terminology into the language of measure-theoretic probability.
We refer to Mohri et al. (2018) [12] as a standard textbook on statistical machine learning theory, and Wainwright (2019) [11] for more mathematical details on the Rademacher complexity and concentration inequalities used in generalization error analysis.
Figure˜1 summarizes the framework of machine learning. Generalization error of a learning algorithm $A:\mathcal{X}^{n}\to\mathcal{F}$ refers to the gap between the original data source (concept $c\in\mathcal{C}$) and the estimated data generator (hypothesis $f\in\mathcal{F}$), which is an output of the learning algorithm $A$ given a dataset ${\bm{x}}\in\mathcal{X}^{n}$ generated according to concept $c$. Following the convention of graphical models, filled-in circles are observable, while unfilled circles are unobservable.
Figure 1: Machine learning framework
2.1 Measure-Theoretic Probability
A probability space $(\Omega,\mathcal{M},\mu)$ is a triple composed of measurable space $\Omega$, $\sigma$-algebra $\mathcal{M}$, and finite positive measure $\mu$ satisfying $\mu(\Omega)=1$.
A random variable $X:\Omega\to\mathcal{X}$ is a measurable map from $\Omega$ to a measurable space $\mathcal{X}$. A realization $x\in\mathcal{X}$ of a random variable $X$ is an image $X(\omega)$ of $X$ at a certain element $\omega\in\Omega$. Following the convention, we use the uppercase for random variables (e.g. $X,Y,\ldots$), while the lowercase for their realizations (e.g. $x,y,\ldots$). The distribution (or law) $P$ of a random variable $X$ is the push-forward measure $X_{\sharp}\mu$ of $\mu$ by $X$. We say $X$ is a random sample in $\mathcal{X}$ drawn according to probability distribution $P$, denoted $X\sim P$, when $X$ is a random variable and its distribution is $P$.
We say a sequence $\bm{X}:=\{X_{k}\}_{k=1}^{n}$ of random variables (or a random vector for short) is independent and identically distributed (i.i.d.), denoted $\bm{X}\sim^{iid}P$, when each $X_{k}$ has the same distribution $P$, and all are mutually independent, so as the distribution $P_{\bm{X}}$ of $\bm{X}$ is given by the component-wise product: $P_{\bm{X}}(\bm{X})=P^{n}(\bm{X}):=\prod_{k=1}^{n}P(X_{k})$.
The expectation of a measurable function $f:\mathcal{X}\to\mathbb{R}$, denoted $\mathbb{E}_{X}[f(X)]$ or $P[f(X)]$, is the integration of $f$ with respect to the measure $P$, that is, $\mathbb{E}_{X}[f(X)]:=\int_{\mathcal{X}}f(x)\mathrm{d}P(x)=\int_{\Omega}f\circ X(\omega)\mathrm{d}\mu(\omega)$. By $L^{1}(P)$, we write the Banach space of all $P$-integrable functions $f:\mathcal{X}\to\mathbb{R}$.
Those concepts are formalized in namespace MeasureTheory in Mathlib. We have formalized the probability space $(\Omega,\mathcal{M},\mu)$ and a random vector $\bm{X}=\{X_{k}\}_{k=1}^{n}:\Omega\to\mathcal{X}^{n}$ as follows:
variable : Type} [MeasurableSpace Ω]
variable : Measure Ω} [IsProbabilityMeasure μ]
variable {甲 : Type} [MeasurableSpace 甲]
variable {n : ℕ} {X : Fin n Ω 甲} (hX : k : Fin n, Measurable (X k))
Especially, the expectation of a map f : Ω is simply denoted as μ[f].
High-probability statement
In probability theory, especially in the context of measure concentration, an event $E$ with parameter $\varepsilon>0$ is said to occur with high probability (w.h.p.) when for every $\varepsilon>0$ there exists $\delta>0$ such that the event $E(\varepsilon)$ occurs with probability at least $1-\delta$. Precisely, it means the following inequality (a.k.a. concentration of probability measure, or tail probability bound, with rate function $\beta$) holds for complement event $E^{c}(\varepsilon)$:
$P(E^{c}(\varepsilon))\leq\beta(\varepsilon),$
or equivalently, for event $E(\varepsilon)$,
$P(E(\varepsilon))\geq 1-\beta(\varepsilon).$
2.2 Statistical Machine Learning
The sample space $\mathcal{X}$ is a measurable set of datasets. Following convention, an observation (an example, or a datum) refers to a single element $x\in\mathcal{X}$, while a dataset (a sample, or a data) refers to a single sequence of observations ${\bm{x}}=\{x_{k}\}_{k=1}^{n}\subset\mathcal{X}^{n}$. For example, in the image recognition problem, a single dataset ${\bm{x}}\in\mathcal{X}^{n}$ is an $n$-fold pairs $\{(\mathtt{image}_{k},\mathtt{label}_{k})\}_{k=1}^{n}$ of images and labels. In statistical machine learning, a single dataset ${\bm{x}}=\{x_{k}\}_{k=1}^{n}$ is formulated as a realization of a random vector $\bm{X}=\{X_{k}\}_{k=1}^{n}:\Omega\to\mathcal{X}^{n}$. If nor necessary, we omit emphasizing the dependency in sample size $n$, and simply write ${\bm{x}},\bm{X},\mathcal{X}^{n}$ as $x,X,\mathcal{X}$.
The concept class $\mathcal{C}$ is a collection of data sources (called concepts) that describe how datasets are obtained. In this study, we assume $\mathcal{C}$ to be a family of random vectors $\bm{X}:\Omega\to\mathcal{X}^{n}$, or equivalently, probability distributions $P$ on $\mathcal{X}^{n}$. In the example of the image recognition problem, a single concept is a random vector $\bm{X}:\Omega\to\mathcal{X}^{n}$, or its law $P_{\bm{X}}(\{\mathtt{image}_{k},\mathtt{label}_{k}\}_{k=1}^{n})$. In statistical machine learning, the concept itself is supposed to be unobservable, and only the dataset to be observable. In other words, only an image ${\bm{x}}=\bm{X}(\omega_{0})$ (at a certain $\omega_{0}\in\Omega$) is given, but the map $\bm{X}:\Omega\to\mathcal{X}^{n}$ itself is not given.
The hypothesis class $\mathcal{F}$ is another collection of data generators (called hypotheses, or learning machines). Like concepts, hypotheses describe the data generation process. However, unlike concepts, hypotheses have parameters, say $\theta\in\Theta$, that we can freely manipulate. For example, in deep learning, the hypothesis class is a set of deep neural networks (DNNs).
A learning algorithm $A$ is a measurable map $\mathcal{X}^{n}\to\mathcal{F}$ that describes how to associate datasets with hypotheses. Regarding the data generation process by concepts as a forward process, learning algorithm $A$ corresponds to a backward process. In the terms of statistical estimation theory, $A$ is an estimator, and the learned machine $f$ is an image $f=A({\bm{x}})$ of a given dataset ${\bm{x}}\in\mathcal{X}^{n}$. For example, in deep learning, $A$ is the process of empirical risk minimization by using stochastic gradient descent on the parameter space of DNNs.
As illustrated in Figure˜1, the generalization error (explained in the next subsection) estimates the discrepancy between the learned hypothesis $f=A({\bm{x}})$ and the original concept $c$. A learning algorithm is considered better if its generalization error is smaller.
We have formalized the hypotheses class $\mathcal{F}=\{f_{i}:\mathcal{X}\to\mathbb{R}\mid i\in\iota\}$ as follows:
variable : Type} [TopologicalSpace ι] [SeparableSpace ι] [FirstCountableTopology ι]
variable {f: ι ℝ} (hf : i, Measurable (f i))
The separable and first-countable assumptions on indeces $\iota$ is required in the formalization of Rademacher complexity, one of the principal terms in the main theorem. These assumptions are rarely made explicit in standard textbooks, and clarified through our formalization.
2.3 Generalization Error Analysis
The training data(set) refers to a random vector $\bm{X}:\Omega\to\mathcal{X}^{n}$, and a test data(set) refers to another random variable $X^{\prime}:\Omega\to\mathcal{X}$ that is statistically independent from training data $\bm{X}$. In this study, we assume (1) that the training dataset is i.i.d., so $\bm{X}\sim^{iid}P$, and (2) that both training and test datasets have the common distribution $P$, so $X^{\prime}\sim P$. We note that $P$ itself is unknown, although we are supposed to know that $\bm{X}$ and $X^{\prime}$ have the same distribution. This may sound technical, but natural when i.i.d. sampling is easy, thus often assumed in the basic setting.
A (pointwise) loss function $\ell:\mathcal{F}\times\mathcal{X}\to\mathbb{R}_{\geq 0}$ is a measurable functional that associates hypotheses with positive numbers. For example, the squared error loss $\ell(f,(x,y)):=|f(x)-y|^{2}$ is typical in supervised learning.
The training error (a.k.a. empirical risk) $L(f\mid\bm{X})$ of a hypothesis $f$ over a training dataset $\bm{X}$ is the sample average of the pointwise loss function: $L(f\mid\bm{X}):=\frac{1}{n}\sum_{k=1}^{n}\ell(f,X_{k})$. We note that in the real-world application, we can only compute its realization, say $L(f\mid{\bm{x}})$.
The test error (a.k.a. population risk) $L(f)$ of a hypothesis $f$ (over a test dataset $X$) is the expectation of the pointwise loss function: $L(f):=\mathbb{E}_{X}[\ell(f,X)]$. Following the convention, the dependency on $X$ is omitted for simplicity. By the assumptions that both training and test datasets are i.i.d. samples, namely $(\bm{X},X)\sim^{iid}P$, the expectation of the training error over the i.i.d. draw of training dataset is identical to the test error: $\mathbb{E}_{\bm{X}}[L(f\mid\bm{X})]=L(f)$.
The test error $L(f)$ is understood as measuring the generalization performance of a hypothesis on unseen data, because the distribution $P$ of test data is unknown. Indeed, it is the (first) definition of the generalization error as explained soon below.
The generalization error refers to three related quantities: (1) population risk (or test error) $L(f)$ itself, (2) generalization gap (the gap between test and training errors) $\Delta(f):=L(f)-L(f\mid\bm{X})$, and (3) excess risk (the population risk relative to its infimum) $L(f)-\inf_{f\in\mathcal{F}}L(f)$. In all three definitions, the interest lies in the (either absolute or relative) value of population risk $L(f)$, and either one will be obtained depending on the estimation technique employed.
In this study, the main theorem (Theorem˜1) presents an upper bound on generalization gap $\Delta(f)$ by using the Rademacher complexity, which estimates the second meaning of generalization error. We note that, as clarified in the remark (Remark˜3), an upper bound of the gap $L(f)-L(f\mid\bm{X})\leq B$ can be trivially turned into the upper bound of the risk $L(f)\leq L(f\mid\bm{X})+B$, which estimates the first meaning of generalization error.
3 Main Results
3.1 Rademacher Complexity
We define the empirical and population Rademacher complexities of a hypothesis class $\mathcal{F}$.
Definition 1 (Rademacher Variable).
A uniform random variable $\sigma$ taking values in $\{\pm 1\}$ is called a Rademacher variable, and an i.i.d. sequence of Rademacher variables ${\bm{\sigma}}:=\{\sigma_{k}\}_{k=1}^{n}$ (i.e. uniform random vector taking values in $\{\pm 1\}^{n}$) is called a Rademacher vector.
Definition 2 (Rademacher Complexity).
Let $\mathcal{F}\subset L^{1}(P)$ be a separable subspace of real-valued integrable functions on $\mathcal{X}$. Let $\bm{X}=\{X_{k}\}_{k=1}^{n}$ be an i.i.d. random vector drawn from distribution $P$. The Empirical Rademacher complexity is defined as
$\mathcal{R}(\mathcal{F}\mid\bm{X})$ $:=\mathbb{E}_{{\bm{\sigma}}}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}\sigma_{k}f(X_{k})\Bigg{|}\right]$
$=\frac{1}{2^{n}}\sum_{{\bm{\sigma}}\in\{\pm 1\}^{n}}\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}\sigma_{k}f(X_{k})\Bigg{|},$
and the (population) Rademacher complexity is defined as its expectation
$\mathcal{R}_{n}(\mathcal{F})$ $:=\mathbb{E}_{\bm{X}}[\mathcal{R}(\mathcal{F}\mid\bm{X})]$
$=\int_{\Omega^{n}}\left[\frac{1}{2^{n}}\sum_{{\bm{\sigma}}\in\{\pm 1\}^{n}}\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}\sigma_{k}f(X_{k}\circ\omega)\Bigg{|}\right]\mathrm{d}\mu^{n}(\omega).$
In our formalization, both the Rademacher vector and empirical Rademacher complexity are formalized without probability for simplicity as follows:
def Signs (n : ℕ) : Type :=
Fin n ({-1, 1} : Finset ℤ)
noncomputable
def empiricalRademacherComplexity
(n : ℕ) (f : ι→甲→ℝ) (x : Fin n→甲) : :=
(Fintype.card (Signs n) : ℝ)⁻¹ *
σ : Signs n, i, |(n : ℝ)⁻¹ * k : Fin n, k : ℝ) * f i (x k)|
On the other hand, the (population) Rademacher complexity is formalized as follows:
noncomputable
def rademacherComplexity
(n:ℕ) (f:ι→甲→ℝ) (μ:Measure Ω) (X:Ω→甲) : :=
μⁿ[fun ω : Fin n Ω empiricalRademacherComplexity n f (X ω)]
Namely, we turn the (non-probabilistic) empirical Rademacher complexity $\mathcal{R}(\mathcal{F}\mid\bm{X})$ into its probabilistic counter by pulling it back to $\Omega$ as $\mathcal{R}(\mathcal{F}\mid\bm{X}\circ\omega)$. Here $\mathtt{\mu^{n}}$ is a local notation for the product measure $\mu^{n}$ defined as follows:
local notation "μⁿ" => Measure.pi (fun _ μ)
Remark 1.
The keyword noncomputable indicates that it invokes a non-constructive operation—taking the supremum (least upper bound) i of a set of real numbers. This operation is provided in mathlib and is implemented using the choice operator Classical.choose. In general, Lean requires any definition that uses Classical.choose to be marked non computable. We emphasize that it is purely a technical annotation demanded by Lean and has no bearing on the mathematical content.
3.2 Generalization Error Bounds by Rademacher Complexity
Here, we explain the main theorem and its formalization. We refer to Theorem 4.10 from Wainwright (2019) [11] and Theorem 3.3 from Mohri et al. (2018) [12]. To be precise, Mohri et al.’s statement is a corollary of Wainwright’s statement tailored for practical purpose. So, we explain Wainwright’s version as the main theorem, and Mohri et al.’s version in the remark.
Theorem 1 (Generalization Error Bound by Rademacher Complexity).
Suppose that the hypothesis class $\mathcal{F}$ (includes the loss function and) is $b$-uniformly bounded, namely there exists a scalar $b\geq 0$ such that $\sup_{f\in\mathcal{F}}\|f\|_{L^{\infty}(\mathcal{X})}\leq b$. For any positive integer $n\geq 1$ and scalar $\varepsilon\geq 0$, the following holds with probability at least $1-\exp\left(-\frac{n\varepsilon^{2}}{2b^{2}}\right)$:
$\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}f(X_{k})-\mathbb{E}_{X}[f(X)]\Bigg{|}\leq 2\mathcal{R}_{n}(\mathcal{F})+\varepsilon$
Following the convention of high-probability statement, the main theorem is formalized as follows:
theorem main
[MeasurableSpace 甲][Nonempty 甲]
[Nonempty ι][TopologicalSpace ι][SeparableSpace ι][FirstCountableTopology ι]
[IsProbabilityMeasure μ]
(f : ι ℝ) (hf : i, Measurable (f i))
(X : Ω 甲) (hX : Measurable X)
{b : ℝ} (hb : 0 b) (hf : i x, |f i x| b)
{t : ℝ} (ht : 0 t) (ht : t * b ^ 2 1 / 2)
: ℝ} (hε : 0 ε) :
(μⁿ (fun ω 2 rademacherComplexity n f μ X + ε uniformDeviation n f μ X (X ω))).toReal
(- ε ^ 2 * t * n).exp := by
Remark 2.
In the main theorem, we assume that a hypothesis includes a loss function. Namely, in the example of image recognition, a hypothesis is not a predictor $g:\mathcal{X}_{\mathtt{image}}\times\Theta\to\mathcal{X}_{\mathtt{label}}$ alone, but a composite $f:\mathcal{X}\times\Theta\to\mathbb{R}_{\geq 0}$, $f((\mathtt{image},\mathtt{label}),\theta):=\ell(g(\mathtt{image},\theta),\mathtt{label})$ of $g$ followed by a loss function $\ell$ such as a cross-entropy or a squared error loss.
Remark 3.
In practice, the quantity of primary interest is the population risk $\mathbb{E}_{X}[\widehat{f}(X)]$ of the hypothesis $\widehat{f}=A(\bm{X})$ obtained by learning algorithm $A$. Because we are only given the training dataset $\bm{X}(\omega)\in\mathcal{X}$ (as a realization), and we do not know the data distribution $P_{\bm{X}}$ itself, this expectation is intractable. Nonetheless, as a consequence of the main theorem, the population risk can be estimated in a tractable manner as follows
$\mathbb{E}_{X}[\widehat{f}(X)]\leq\frac{1}{n}\sum_{k=1}^{n}\widehat{f}(X_{k})+2\mathcal{R}_{n}(\operatorname{\mathrm{Im}}A)+\sqrt{\frac{2b^{2}\log 1/\delta}{n}}$
with probability at least $1-\delta$ over the draw of an i.i.d. sample $\bm{X}$. This is Mohri et al’s version of the main theorem (Theorem 3.3 in [12]). We note that we further need to compute the Rademacher complexity separately depending on the specific problem.
Proof.
The proof is two-fold: Use McDiarmid’s (bounded difference) inequality, and symmetrization argument. Put the supremum $\Delta(\mathcal{F}\mid\bm{X})$ of the absolute deviation and its expectation $\Delta(\mathcal{F})$ as follows:
$\Delta(\mathcal{F}\mid\bm{X})$ $:=\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}f(X_{k})-\mathbb{E}_{X}[f(X)]\Bigg{|},$
$\Delta(\mathcal{F})$ $:=\mathbb{E}_{\bm{X}}[\Delta(\mathcal{F}\mid\bm{X})].$
We call $\Delta(\mathcal{F}\mid\bm{X})$ the uniform deviation for short, and formalize it as follows:
def uniformDeviation
(n : ℕ) (f: ι→甲→ℝ) : Measure Ω)
(X : Ω 甲) : (Fin n 甲) :=
fun y i, |(n : ℝ)⁻¹ * k : Fin n, f i (y k) - μ [fun ω’ f i (X ω’) ]|
By McDiarmid’s inequality, the deviation of the uniform deviation from its mean is upper bounded by $\varepsilon$ as
$\Delta(\mathcal{F}\mid\bm{X})-\Delta(\mathcal{F})<\varepsilon$ (1)
with probability at least $1-\exp\left(-\frac{n\varepsilon^{2}}{2b^{2}}\right)$. In other words, the following inequality holds:
$\mu\left\{\omega\mid\Delta(\mathcal{F}\mid\bm{X})(\omega)-\Delta(\mathcal{F})\geq\varepsilon\right\}\leq\exp\left(-\frac{n\varepsilon^{2}}{2b^{2}}\right),$
which is formalized as follows:
theorem uniformDeviation_mcdiarmid
[MeasurableSpace 甲][Nonempty 甲]
[Nonempty ι][TopologicalSpace ι][SeparableSpace ι][FirstCountableTopology ι]
[IsProbabilityMeasure μ]
{X : Ω→甲} (hX : Measurable X)
(hf : i, Measurable (f i))
{b : ℝ} (hb : 0 b) (hf’: i x, |f i x| b)
{t : ℝ} (ht : 0 t) (ht : t * b ^ 2 1 / 2)
: ℝ} (hε : 0 ε) :
(μⁿ (fun ω : Fin n Ω uniformDeviation n f μ X (X ω) -
μⁿ[fun ω : Fin n Ω uniformDeviation n f μ X (X ω)] ε)).toReal
(- ε ^ 2 * t * n).exp := by
Based on the following symmetrization argument, $\Delta(\mathcal{F})$ is estimated by the Rademacher complexity as follows. Take another i.i.d. sequence $\bm{Y}:=\{Y_{k}\}_{k=1}^{n}\sim^{iid}P$ independent of $\bm{X}$. Then,
$\Delta(\mathcal{F})$ $=\mathbb{E}_{\bm{X}}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}\left\{f(X_{k})-\mathbb{E}_{Y_{k}}[f(Y_{k})]\right\}\Bigg{|}\right]$
$=\mathbb{E}_{\bm{X}}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\mathbb{E}_{\bm{Y}}\left[\frac{1}{n}\sum_{k=1}^{n}\left\{f(X_{k})-f(Y_{k})\right\}\right]\Bigg{|}\right]$
$\leq\mathbb{E}_{\bm{X},\bm{Y}}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}\left\{f(X_{k})-f(Y_{k})\right\}\Bigg{|}\right]$
$=\mathbb{E}_{\bm{X},\bm{Y},{\bm{\sigma}}}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}\sigma_{i}\left\{f(X_{k})-f(Y_{k})\right\}\Bigg{|}\right]$
$\leq 2\mathbb{E}_{\bm{X},{\bm{\sigma}}}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}\sigma_{i}f(X_{k})\Bigg{|}\right]=2\mathcal{R}_{n}(\mathcal{F}).$ (2)
This argument is formalized as the proof term of the following theorem:
theorem le_two_smul_rademacher
[Nonempty ι][TopologicalSpace ι][SeparableSpace ι]
[IsProbabilityMeasure μ]
(X : Ω 甲) (hf : i, Measurable (f i X))
{b : ℝ} (hb : 0 b) (hf’: i x, |f i x| b) :
μⁿ[fun ω : Fin n Ω uniformDeviation n f μ X (X ω)]
2 rademacherComplexity n f μ X := by
Finally, the combination of the estimates (1) and (2) yields the assertion. ∎
3.3 McDiarmid’s Inequality
McDiarmid’s inequality, a.k.a. the bounded difference inequality, estimates the concentration bound of a function $f:\mathcal{X}^{n}\to\mathbb{R}$ around its mean $\mathbb{E}[f(\bm{X})]$ under the assumption that $f$ satisfies the bounded difference property. We note that the case when $f$ is a sum $f(\bm{X})=\sum_{k=1}^{n}X_{k}$ reproduces Hoeffding’s inequality. We refer to Corollary 2.21 from [11] for more details.
There are two ways to prove this inequality: directly by using Hoeffding’s lemma (explained later), or indirectly as a corollary of the Azuma-Hoeffding inequality (see e.g. Corollary 2.20 in [11]). In this study, we employed the former direct way. We remark that the Azuma-Hoeffding inequality was not been formalized when we were developing the formalization, but now it is formalized in Mathlib.Probability.Moments. SubGaussian.
Definition 3 (Bounded Difference Property).
Given an $n$-tuple ${\bm{x}}\in\mathcal{X}^{n}$ and an element $x^{\prime}\in\mathcal{X}$, let ${\bm{x}}(k,x^{\prime})$ denote a new $n$-tuple obtained by replacing the $k$-th component $x_{k}$ of $\bm{x}$} with $x^{\prime}$. A function $f:\mathcal{X}^{n}\to\mathbb{R}$ satisfies the bounded difference property if there exists a sequence $\{c_{k}\}_{k=1}^{n}$ of positive numbers such that for all $k\in[n],{\bm{x}}\in\mathcal{X}^{n},x^{\prime}\in\mathcal{X}$:
$|f({\bm{x}}(k,x^{\prime}))-f({\bm{x}})|\leq c_{k}.$
Theorem 2 (One-sided McDiarmid’s Inequality, or Bounded Differences -).
Suppose that a measurable function $f:\mathcal{X}^{n}\to\mathbb{R}$ satisfies the bounded difference property with bounds $\{c_{k}\}_{k=1}^{n}$, and suppose that a real number $t\in\mathbb{R}$ satisfies $t\leq 1/\sum_{k=1}^{n}c_{k}^{2}$. Let $\bm{X}=\{X_{k}\}_{k=1}^{n}\sim P$ be an i.i.d. sequence. Then, for any $\varepsilon\geq 0$, we have:
$\mu\left\{\omega\mid f(\bm{X})(\omega)-\mathbb{E}_{\bm{X}}[f(\bm{X})]\geq\varepsilon\right\}\leq\exp\left(-2\varepsilon^{2}t\right)$
theorem mcdiarmid_inequality_pos
(X : ι→Ω→甲) (hX : i, Measurable (X i)) (hX : iIndepFun X μ)
(c : ι→ℝ)
(f : (ι→甲)→ℝ) (hf : Measurable f) (hf : (i : ι) (x : ι 甲) (x : 甲),
|f x - f (Function.update x i x’)| c i) bounded difference property
: ℝ) (hε : ε>0)
(t : ℝ) (ht : t * i, (c i) ^ 2 1) :
(fun ω : Ω (f (Function.swap X)) ω - μ[f (Function.swap X)] ε)).toReal
(-2 * ε ^ 2 * t).exp := by
3.4 Hoeffding’s Lemma
Hoeffding’s lemma states that an almost surely bounded random variable $X$ is sub-Gaussian. It is used to show Hoeffding’s inequality and its generalization McDiarmid’s inequality. We refer to Lemma D.1 in [12] for more details. In the proof, we use exponential tilting, which has already been implemented in Mathlib.Probability.Moments.Tilted.
Theorem 3 (Hoeffding’s Lemma).
For a real random variable $X$ with $\mathbb{E}[X]=0$ and $X\in[a,b]$ almost surely, the inequality
$\mathbb{E}_{X}\left[\exp tX\right]\leq\exp\left(\frac{t^{2}(b-a)^{2}}{8}\right)$
holds almost surely for all $t\in\mathbb{R}$.
theorem hoeffding
[IsProbabilityMeasure μ]
(t a b : ℝ) {X : Ω ℝ} (hX : AEMeasurable X μ)
(h : ∀ᵐ ω ∂μ, X ω Set.Icc a b) (h0 : μ[X] = 0) :
mgf X μ t exp (t^2 * (b - a)^2 / 8) := by
Here, mgf is the moment generating function $\mathbb{E}_{X}[\exp tX]$ of a real random variable $X$ defined in Mathlib.Probability. Moments.Basic as follows:
def mgf (X : Ω ℝ) : Measure Ω) (t : ℝ) : :=
μ[fun ω => exp (t * X ω)]
4 Example: $L^{2}$-Regularized Regression
As an application, we present the generalization error bound of $L^{2}$-regularized linear regression models. Let $\langle\bullet,\bullet\rangle_{2}$ and $|\bullet|_{2}$ denote the $d$-dimensional dot product and its induced norm, respectively. Suppose that both input space $\mathcal{X}$ and parameter space $\mathcal{W}$ are closed $\ell^{2}$-balls: $\mathcal{X}=\{{\bm{x}}\in\mathbb{R}^{d}\mid|{\bm{x}}|_{2}\leq B_{X}\}$ and $\mathcal{W}=\{{\bm{w}}\in\mathbb{R}^{d}\mid|{\bm{w}}|_{2}\leq B_{W}\}$, so that the hypotheses class $\mathcal{F}$ is the collection of linear regression models with bounded parameters: $\mathcal{F}=\{{\bm{x}}\mapsto\langle{\bm{w}},{\bm{x}}\rangle_{2}\mid{\bm{w}}\in\mathcal{W}\}$. Then, the empirical Rademacher complexity of $\mathcal{F}$ is bounded as follows:
$\mathcal{R}(\mathcal{F}\mid\bm{X})=\mathbb{E}_{{\bm{\sigma}}}\left[\sup_{|{\bm{w}}|_{2}\leq B_{W}}\Bigg{|}\frac{1}{n}\sum_{k=1}^{n}\sigma_{k}\langle{\bm{w}},X_{k}\rangle_{2}\Bigg{|}\right]\leq\frac{B_{W}B_{X}}{\sqrt{n}}.$
theorem linear_predictor_l2_bound
[Nonempty ι](d:ℕ)(bW bX : ℝ)(hx : 0≤bX)(hw : 0≤bW)
(X : Fin n Metric.closedBall (0 : EuclideanSpace (Fin d)) bX)
(W : ι Metric.closedBall (0 : EuclideanSpace (Fin d)) bW) :
empiricalRademacherComplexity n
(fun (i : ι) x => ⟪((Subtype.val W) i), x’⟫) (Subtype.val X)
bX * bW / √(n : ℝ) := by
Here $\langle\!\langle\bullet,\bullet\rangle\!\rangle$ is a local notation for the dot product defined as follows:
local notation "⟪" x ", " y "⟫" => @inner _ _ x y
5 Behind-the-Scenes Stories
We review the particularly challenging components of the development and the methodological choices that enabled progress.
5.1 Independent Variables
In the Rademacher complexity, the training dataset is assumed to be i.i.d.. This is essential, for example, in the symmetric arguments. In a textbook-style formulation, i.i.d. variables $X_{1},\ldots,X_{n}$ are formulated as a map $\mathrm{Fin}\,n\to\Omega\to\mathcal{X}$. However, we noticed that it is much convenient to formalize them as compositions of a single variable $X:\Omega\to\mathbb{R}$ with the coordinate projections $\Omega^{n}\to\Omega$, letting $\Omega^{n}$ as the base probability space. In this view, independence need not be explicitly assumed as it follows directly from the construction, which helps simplifying descriptions of theorems.
On the other hand, in the McDiarmid’s inequality, the random variables $X_{1},\ldots,X_{n}$ are assumed to be independent but need not be identically distributed. So we formalize them as distinct functions $X:\mathrm{Fin}\,n\to\Omega\to\mathbb{R}$ and explicitly assume independence as an explicit condition. Here the base probability space is $\Omega$.
The fact that the former construction $\Omega^{n}\to\Omega\to\mathcal{X}$ satisfies the independence condition required in the latter is a theorem that should be proved. We could not find this result in Mathlib, so we supplied our own proof. We also attempted another formalization that defines Rademacher complexity via probability mass functions, but deriving its properties along this path proved difficult, and we abandoned that approach.
5.2 Topological Details of Index Set $\iota$ and Hypotheses Class $\mathcal{F}$
The Rademacher complexity is an expectation of the supremum over hypotheses class $\mathcal{F}$, and in modern machine learning settings, $\mathcal{F}$ is often uncountable. When $\mathcal{F}$ is uncountable, however, measurability is not generally preserved under pointwise $\sup$. We therefore first prove results in the countable case and then extend to the separable case. Even when uncountable, a separable family allows the $\sup$ of continuous functions to be computed over a countable dense subset, reducing to the countable case. However, this requires that continuity be preserved under integration, which in turn requires first countability (a point that is, to our knowledge, not emphasized in standard textbooks). These conditions are satisfied in the spaces relevant to our applications, so they pose no practical obstacle.
5.3 Conditional Expectation
Textbook proofs of McDiarmid’s inequality typically proceed via conditional expectations. However, because conditional expectation is defined abstractly, it is difficult to carry out concrete computations directly from that definition. (At the time we were developing the formalization, Mathlib have not yet included the Doob–Dynkin lemma; now there is Mathlib.MeasureTheory.Function.FactorsThrough). We therefore avoided conditional expectations and instead defined the relevant quantities directly by integration, using independence. This changes the proof order: in the original argument the constructed sequence $Y$ is a martingale by construction, whereas in our approach (without conditional expectations) we establish the martingale property from independence.
5.4 Integral and Supremum
A large fraction of the lines in many lemmas is devoted to handling integrability conditions (and the treatment of $\sup$), which are often considered routine. Each time we performed an algebraic manipulation under the integral sign, we had to supply a fresh proof of integrability.
6 Literature Overview
Tables˜1 and 2 summarize close literature to this study.
6.1 PAC Learning and VC Dimension (Lean, Rocq)
Table 1: Formalization of Machine Learning Theory
Bentkamp et al. (2016,2019) [5, 6] Isabelle/HOL Expressive power superiority of deep over shallow
Bagnall and Stewart (2019) [3] Coq Generalization error bounds for finite hypothesis class
Tassarotti et al. (2021) [13] Lean PAC learnability of decision stumps
Vajjha et al. (2021) [16] Coq Convergence of reinforcement learning algorithms
Vajjha et al. (2022) [17] Coq Stochastic approximation theorem
Hirata (2025) [9] Isabelle/HOL No free lunch theorem
(ours) Lean Generalization error bound by Rademacher complexity
Table 2: Formalization of Concentration Inequalities
Markov/Chebyshev Lean (Mathlib), Coq (MathComp-Analysis[1], IBM/FormalML[16]),
Isabelle/HOL (HOL-Probability)
Azuma-Hoeffding Lean (Mathlib)
McDiarmid (ours), Isabelle/HOL (AFP) [10]
A hypothesis class with finite VC dimension (or simply a finite class) admits Probably Approximately Correct (PAC) generalization bounds. Early formalizations focused on specific cases as follows:
Finite Hypothesis Classes:
Bagnall & Stewart (2019) [3] proved a general PAC bound in Rocq for any finite hypothesis class using Hoeffding’s inequality. Essentially, if $\mathcal{F}$ is finite, with probability $1-\delta$ the true error is within $\varepsilon$ of the training error for $n\gtrsim\frac{1}{\varepsilon^{2}}(\ln\|\mathcal{F}\|+\ln\frac{1}{\delta})$. Their Rocq development (part of the MLCERT system) used a union bound and Chernoff/Hoeffding bounds to link training and test errors. This formal result was applied to certify small neural network models’ performance. However, it was limited to finite $\mathcal{F}$ (not covering infinitely large model classes).
Decision Stump Class ($VC=1$):
Tassarotti et al. (2021) [13] gave a full Lean 3 proof that the concept class of decision stumps (threshold classifiers in $\mathbb{R}$) is PAC-learnable. This is a classic textbook example with VC dimension 1. The formal proof uncovered subtle measure-theoretic issues that are glossed over in informal proofs. For instance, textbooks often assume measurability of argmax operations on sample data without proof—the formalization had to rigorously prove measurability and proper probability space definitions for the learning algorithm. The authors structured the proof to separate combinatorial reasoning about the algorithm’s behavior from the analytic reasoning about probabilities. They employed the Giry monad (in Lean’s category theory library) to handle distributions, and ultimately derived the standard PAC guarantee for decision stumps.
6.2 Formalizing Machine Learning Theory
Statistical learning theory extends beyond classical generalization bounds.
Expressiveness of Deep Neural Networks:
Bentkamp et al. (2016,2019) [5, 6] formalized in Isabelle/HOL a theorem that deep networks can represent certain functions exponentially more efficiently than shallow ones. The formalization, simplified and generalized the original proof in 2016 within Isabelle’s logic. To support the proof, Bentkamp developed libraries for linear algebra (matrix ranks), multivariate polynomials, and even Lebesgue measure integration. The result is not about generalization error, but rather about the capacity/representation power of deep vs. shallow networks—nevertheless, it showcases the application of proof assistants to core theoretical ML questions. It also enriched Isabelle’s libraries with notions like tensor products and rank, which are useful in learning theory.
Convergence of Reinforcement Learning:
Vajjha et al. (2021) [16] verified the convergence of value iteration and policy iteration for Markov Decision Processes in Rocq. In a follow-up, they formalized a stochastic approximation theorem [17] useful for analyzing RL algorithms. These efforts, under the IBM FormalML project, build a bridge between learning theory and formal verification by proving probabilistic convergence properties in Rocq. They required heavy use of Rocq’s analysis libraries and bespoke techniques (e.g. coinduction for probabilistic processes).
6.3 Concentration Inequalities and Background Formalizations
Many learning-theoretic proofs rely on concentration of measure results and related probabilistic inequalities. Over the 2010s–2020s, these foundational results have been increasingly formalized, often as prerequisites for the theorems above:
Isabelle/HOL:
Its probability theory library (HOL-Probability) already included basic inequalities like Markov’s inequality, Chebyshev’s inequality, and exponential tail (Chernoff/Hoeffding) bounds by the late 2010s. In 2023, Karayel and Tan [10] contributed an AFP entry “Concentration Inequalities” which adds more advanced results. This includes Bennett’s and Bernstein’s inequalities (for sub-exponential random variables), Efron–Stein’s inequality (variance bound via variance decomposition), McDiarmid’s inequality (bounded differences), and the Paley–Zygmund inequality. Thanks to this, Isabelle/HOL now boasts one of the most extensive collections of concentration results, all formally proven. For example, the formal McDiarmid inequality in Isabelle was crucially used as a point of comparison for the Lean development.
Lean
Lean’s mathlib gained basic measure theory around 2019, including Lebesgue integration and independence. By mid-2020s, mathlib had formal proofs of Markov and Chebyshev inequalities. This study also introduced Hoeffding’s lemma (a result that a bounded zero-mean variable is sub-Gaussian) and then proved Hoeffding’s inequality as a corollary of Chernoff bound techniques. It also implemented McDiarmid’s inequality from scratch in Lean 4. Azuma-Hoeffding was implemented independently around the same time with this study.
Rocq
Historically, Rocq’s standard library did not include measure-theoretic concentration results. But developments like MathComp-Analysis [1] and research projects have added some pieces. Affeldt and others built a formal measure theory in Rocq compatible with the Mathematical Components library. Using these, the IBM FormalML team and others have formalized at least the basic inequalities (Markov, Chebyshev) in Rocq. Hoeffding’s and McDiarmid’s inequalities were not present in Rocq as of early 2020s except in special-case proofs (e.g. MLCert implicitly used a form of Hoeffding’s bound for finite samples).
7 Conclusion
In this study, we formalized the Rademacher complexity to bound the generalization error for the first time in Lean 4. The formal proof mirrors textbook treatments (e.g. Mohri et al., 2018; Wainwright, 2019) but with all measurability and integration details rigorously certified. It lays a foundation to formally verify generalization guarantees in modern settings.
A related advanced result is Dudley’s entropy integral bound (a sharp bound via covering numbers and chaining). As of 2025, no full formal proof of Dudley’s theorem is reported – formalizing it would require developing theory of covering numbers and chaining, which remains an open challenge.
Acknowledgments
We would like to thank the mathlib community members, particularly Rémy Degenne and Yaël Dillies, for their valuable comments regarding formalization of Hoeffding’s lemma. We would like to thank Reynald Affeldt, Alessandro Bruni, Tetsuya Sato, Hiroshi Unno, and Yoshihiro Mizoguchi for having productive discussion on our formalization project. This work was supported by JSPS KAKENHI 24K21316, and JST Moonshot R&D Program JPMJMS2236, PRESTO JPMJPR2125, and BOOST JPMJBY24E2.