### v4

# Lean Formalization of Generalization Error Bound by Rademacher Complexity and Dudley’s Entropy Integral

Abstract
Understanding and certifying the generalization performance of machine learning algorithms—i.e. obtaining theoretical estimates of the test error from a finite training sample—is a central theme of statistical learning theory. Among the many complexity measures used to derive such guarantees, Rademacher complexity yields sharp, data-dependent bounds that apply well beyond classical 011 classification. In this study, we formalize the generalization error bound by Rademacher complexity in Lean 4, building on measure-theoretic probability theory available in the Mathlib library. Our development provides a mechanically-checked pipeline from the definitions of empirical and expected Rademacher complexity, through a formal symmetrization argument and a bounded-differences analysis, to high-probability uniform deviation bounds via a formally proved McDiarmid inequality. A key technical contribution is a reusable mechanism for lifting results from countable hypothesis classes (where measurability of suprema is straightforward in Mathlib) to separable topological index sets via a reduction to a countable dense subset. As worked applications of the abstract theorem, we mechanize standard empirical Rademacher bounds for linear predictors under $\ell_{2}$ and $\ell_{1}$ regularization, and we also formalize a Dudley-type entropy integral bound based on covering numbers and a chaining construction. The code is available at https://github.com/auto-res/lean-rademacher.
1 Introduction
Generalization analysis is the task of estimating, from finite data, how accurately a learned model will perform on unseen test data. Formally, let $\mathcal{X}$ be a data domain and let $\mu$ be an unknown data-generating distribution on $\mathcal{X}$. A learning algorithm observes an i.i.d. sample $S=(x_{1},\dots,x_{n})$ drawn from $\mu$ and selects a hypothesis $h$ from some hypothesis class $\mathcal{H}$ (or, more generally, a function from a parameterized family). If $\ell(h,x)$ denotes a loss value on data $x$, the population risk (test error) and the empirical risk (training error) are
$L(h):=\mathbb{E}_{x\sim\mu}[\ell(h,x)],\qquad\widehat{L}_{S}(h):=\frac{1}{n}\sum_{k=1}^{n}\ell(h,x_{k}).$
A typical goal in statistical learning theory is to bound the generalization gap $|\widehat{L}_{S}(h)-L(h)|$, uniformly over all hypotheses in the class (or at least over the hypotheses produced by a learning algorithm), with high probability over the sample draw.
In this study, we formalize the generalization error bound by Rademacher complexity [5] in Lean 4 [8] based on probability theory formalized in Lean 4’s mathematical library, Mathlib 4 [16]. We refer to Mohri et al. (2018) [13] as a standard textbook on statistical machine learning theory, and Wainwright (2019) [12] for more mathematical details on the Rademacher complexity and concentration inequalities used in generalization error analysis. Concretely, rather than fixing a particular learning algorithm, we formalize a uniform law of large numbers (ULLN) style deviation functional: for a family of measurable real-valued functions $F=\{f_{i}:\mathcal{X}\to\mathbb{R}\}_{i\in\iota}$ we consider
$\mathrm{UD}_{n}(F;S):=\sup_{i\in\iota}\left|\frac{1}{n}\sum_{k=1}^{n}f_{i}(x_{k})-\mathbb{E}_{x\sim\mu}[f_{i}(x)]\right|.$
This quantity subsumes classical generalization-gap bounds by taking $f_{i}$ to be loss functions $\ell(h_{i},\cdot)$, and it also aligns well with the measure-theoretic infrastructure in Mathlib. In the Lean code, the corresponding objects are defined as uniformDeviation and bounded using empiricalRademacherComplexity and rademacherComplexity (Section 4 of the main body).
Why Rademacher complexity?
Classical sample-complexity theory for the PAC learning framework [19] often proceeds via the Vapnik–Chervonenkis (VC) dimension [20]. While powerful, VC-dimension-based bounds are typically data-independent and tailored to worst-case uniform guarantees for 011 classification. In modern learning settings, however, one often seeks bounds that (i) apply to real-valued losses and function classes beyond binary classifiers, and (ii) adapt to the observed sample through a data-dependent complexity. Rademacher complexity provides precisely such a tool: it measures the ability of a function class to correlate with random signs, and yields sharp uniform deviation bounds for a wide variety of learning problems, including analyses that underpin kernel methods and deep learning [13, 15, 3]. This flexibility motivates a foundational formalization effort: once the abstract Rademacher pipeline is mechanized, many concrete generalization analyses can be obtained as specializations by instantiating the function class and bounding its complexity.
A textbook-level bound and its formal counterpart.
At a high level, the classical Rademacher approach yields the following pattern. Assume a uniform bound $|f_{i}(x)|\leq b$. Then one first proves an expected bound
$\mathbb{E}[\mathrm{UD}_{n}(F;S)]\leq 2\mathfrak{R}_{n}(F),$
where $\mathfrak{R}_{n}(F)$ is the (distribution-dependent) Rademacher complexity of $F$. Then one upgrades this to a high-probability statement using a concentration inequality (typically McDiarmid’s inequality), obtaining for all $\delta\in(0,1)$
$\mathbb{P}\left(\mathrm{UD}_{n}(F;S)\leq 2\mathfrak{R}_{n}(F)+b\sqrt{\frac{2\log(1/\delta)}{n}}\right)\geq 1-\delta.$
Our Lean development follows and formalizes this standard “symmetrization + bounded differences” route, and exposes the main tail bounds as theorems uniform_deviation_tail_bound_countable_of_pos (countable index type) and uniform_deviation_tail_bound_separable_of_pos (separable index set) in FoML/Main.lean.
Why Lean and Mathlib?
A recurring theme in modern learning theory is that generalization bounds are most naturally expressed and proved in a measure-theoretic language: product measures for i.i.d. sampling, measurability of suprema, integrability of envelopes, and concentration inequalities stated for measurable functions on probability spaces. Over the past several years, measure theory and probability theory in Lean’s ecosystem have developed rapidly, and Mathlib has become a particularly convenient target for such work due to its extensive analysis libraries and the large community of mathematically oriented contributors. These factors make Lean 4 + Mathlib 4 a natural setting for formalizing generalization analysis at a level close to contemporary textbooks (as opposed to highly simplified, finite or discrete proxies).
Novelty and contributions.
Prior work has formalized generalization bounds, but only in relatively simplified settings. For example, Bagnall and Stewart (2019) [4] formalize a generalization bound for neural networks, but restrict parameters to a finite set, which is insufficient for formalizing today’s diverse generalization analyses. Our contribution is, to our knowledge, the first systematic formalization of generalization analysis via Rademacher complexity. More specifically, the repository provides:
  • a reusable Lean formalization of empirical and expected Rademacher complexity and the associated uniform deviation functional;
  • a mechanically checked symmetrization argument connecting uniform deviations to Rademacher averages;
  • a formally verified bounded-differences (Lipschitz) analysis of the uniform deviation functional and a full McDiarmid inequality proof, yielding high-probability generalization bounds;
  • a principled extension from countable index sets to separable topological parameter spaces via a countable dense reduction, resolving measurability issues of uncountable suprema;
  • worked applications: formal generalization-relevant Rademacher bounds for $\ell_{2}$- and $\ell_{1}$-regularized linear predictors, and a formal Dudley entropy integral bound via covering numbers and chaining.
Paper organization.
Section 2 summarizes the repository at a conceptual level. Sections 3–5 introduce the common measure-theoretic setting, define the core objects, and present the main Rademacher-based expectation and tail bounds. Section 6 develops the $\ell_{2}$ and $\ell_{1}$ linear predictor examples. Section 7 presents the Dudley entropy integral formalization. Section 8 discusses key design choices and engineering challenges that arise when translating textbook arguments into Mathlib’s formal framework, and Section 9 situates the contribution in the broader literature on formalized learning theory and concentration inequalities.
2 Summary of Repository
The repository’s core contribution is a mechanically-checked pipeline from (i) definitions of empirical and expected Rademacher complexity, through (ii) a formal symmetrization argument, to (iii) high-probability uniform deviation (ULLN-style) bounds via a formally proved McDiarmid inequality, with an additional reduction step that extends the results from countable hypothesis classes to separable topological index sets. It also includes worked examples showing how $\ell_{2}$- and $\ell_{1}$-type norm constraints yield standard $O(1/\sqrt{n})$ and $O(\sqrt{\log(d)/n})$ empirical Rademacher-complexity bounds for linear prediction, and a formal Dudley entropy-integral upper bound for empirical Rademacher complexity (via covering numbers and a chaining construction).
From a theorem proving perspective, the notable aspects are: the explicit handling of measurability/integrability hypotheses (often forcing Countable index sets, or separability+continuity to reduce to a countable dense subclass), the use of product measures $\mu^{n}$ rather than i.i.d. sequences as primitive, and the systematic bridging between informal learning-theory arguments and Mathlib’s measure-theoretic infrastructure. Conceptually, the main theorems align with classical results in Rademacher complexity and generalization theory as in Bartlett–Mendelson (risk bounds via Rademacher averages), McDiarmid’s bounded-differences method, and Dudley’s metric-entropy/chaining bounds.
3 Repository Scope, and Assumptions
Our Lean development is structured around a general measure-theoretic formulation of empirical processes: a probability space $(\Omega,m\Omega,\mu)$, a data space $\mathcal{X}$ (sometimes Z in code), and a random variable $X:\Omega\to\mathcal{X}$ (encoding the data distribution). Fix an index type $\iota$ (parameterizing hypotheses), and a family of real-valued functions
$\mathcal{F}\;=\;\{f_{i}:\mathcal{X}\to\mathbb{R}\}_{i\in\iota}.$
In the code, such a class is represented by a single curried term f : ι .
Samples of size $n$ are modeled as functions $\omega:\mathrm{Fin}\ n\to\Omega$, pushed forward through $X$ to obtain $S:=X\circ\omega:\mathrm{Fin}\ n\to\mathcal{X}$. Expectations are integrals in Mathlib notation, e.g. μ[g] for $\int g\,d\mu$.
We assume the reader is comfortable with basic probability/analysis (product measures, integrals, concentration inequalities, norms) but not necessarily with ML-specific vocabulary (hypothesis classes, generalization error, empirical risk). We will use natural logarithms $\log$ (Lean’s Real.log) and the convention that $n\in\mathbb{N}$, $n>0$ when dividing by $n$ or using $\sqrt{n}$. We interpret toReal of an ENNReal probability as the ordinary probability in $[0,1]$ (a standard Mathlib bridge).
4 Core Definitions and Notational Conventions
The central definitions in the file FoML/Defs.lean include the empirical and (expected) Rademacher complexities:
4.1 Rademacher signs as a finite type.
The repository models a sign vector as a function Fin n {-1,1} in $\mathbb{Z}$, with Fintype structure giving a finite uniform averaging over all $2^{n}$ sign assignments.
FoML/Defs.lean
def Signs (n : ℕ) : Type := Fin n ({-1, 1} : Finset ℤ)
Mathematically, one can view Signs n as $\{\pm 1\}^{n}$ equipped with the uniform distribution; Lean often uses explicit finite sums multiplied by $(\#\mathrm{Signs}\ n)^{-1}$ instead of an expectation over i.i.d. $\sigma_{k}$.
4.2 Empirical Rademacher complexity.
For a sample $S:\mathrm{Fin}\ n\to\mathcal{X}$ and a class $f:\iota\to\mathcal{X}\to\mathbb{R}$, the code defines:
$\widehat{\mathfrak{R}}_{n}(f;S):=\frac{1}{|\{\pm 1\}^{n}|}\sum_{\sigma\in\{\pm 1\}^{n}}\sup_{i\in\iota}\left|\frac{1}{n}\sum_{k=1}^{n}\sigma_{k}\,f_{i}(S_{k})\right|.$
This matches the common ML definition up to the choice of explicit averaging versus probabilistic expectation.
FoML/Defs.lean
def empiricalRademacherComplexity (n : ℕ) (f : ι ℝ) (S : Fin n 甲) : :=
(Fintype.card (Signs n) : ℝ)⁻¹ *
σ : Signs n, i, |(n : ℝ)⁻¹ * k : Fin n, k : ℝ) * f i (S k)|
4.3 Expected (distribution-dependent) Rademacher complexity.
For a random variable $X:\Omega\to\mathcal{X}$ and probability measure $\mu$ on $\Omega$, the development forms the product measure $\mu^{n}$ on $\Omega^{n}$ and defines:
$\mathfrak{R}_{n}(f;\mu,X):=\mathbb{E}_{\omega\sim\mu^{n}}\big[\widehat{\mathfrak{R}}_{n}(f;X\circ\omega)\big].$
FoML/Defs.lean
def rademacherComplexity (n : ℕ) (f : ι ℝ) : Measure Ω) (X : Ω 甲) : :=
μⁿ[fun ω : Fin n Ω empiricalRademacherComplexity n f (X ω)]
4.4 Uniform deviation (ULLN functional).
The central “generalization” quantity is the supremum of the absolute difference between the empirical mean and the population mean:
$\mathrm{UD}_{n}(f;\mu,X;S):=\sup_{i\in\iota}\left|\frac{1}{n}\sum_{k=1}^{n}f_{i}(S_{k})\;-\;\mathbb{E}_{\mu}[f_{i}(X)]\right|.$
FoML/Defs.lean
def uniformDeviation (n : ℕ) (f : ι ℝ) : Measure Ω) (X : Ω 甲) (S : Fin n 甲) : :=
i, |(n : ℝ)⁻¹ * k : Fin n, f i (S k) - μ[fun ω’ f i (X ω’)]|
4.5 Two technical variants used later.
The repo also defines a without_abs empirical Rademacher complexity, dropping the absolute value inside the supremum. It proves a comparison lemma (empiricalRademacherComplexity_without_abs_le_empirical RademacherComplexity) under boundedness assumptions, used to pass Dudley/Massart-style bounds (often stated without absolute values) to the absolute-value version when needed.
5 Generalization via Rademacher Complexity and Concentration
5.1 Symmetrization as a first-class formal object.
The file FoML/Symmetrization.lean develops an integral identity that replaces a supremum over differences of two independent samples by an average over sign-flipped sums. This is the backbone of standard Rademacher complexity arguments, and it is exposed as a theorem used downstream: abs_symmetrization_equation. The formal version is phrased over the product measure on $(\Omega\times\Omega)^{n}$ and uses finitary sign averaging.
5.2 Expected supremum bound (Wainwright-style ULLN in expectation).
In FoML/Rademacher.lean, the key theorem expectation_le_rademacher states (informally):
$\mathbb{E}_{\omega\sim\mu^{n}}\left[\sup_{i\in\iota}\left|\sum_{k=1}^{n}f_{i}(X(\omega_{k}))-n\,\mathbb{E}_{\mu}[f_{i}(X)]\right|\right]\;\leq\;2n\cdot\mathfrak{R}_{n}(f;\mu,X),$
under Nonempty ι, Countable ι, measurability hypotheses, and a uniform bound $|f_{i}(z)|\leq b$. The proof explicitly composes: (i) rewriting the mean as a coordinatewise mean on the product space, (ii) applying the symmetrization identity, and (iii) splitting the supremum via triangle inequality (a step corresponding to the classical “ghost sample + symmetrization $\Rightarrow 2\mathfrak{R}$” argument). For context, this is the standard route by which Rademacher complexity controls expected uniform deviations and risk bounds.
FoML/Rademacher.lean (excerpt)
theorem expectation_le_rademacher {X : Ω Z} [Nonempty ι] [Countable ι]
(hf : i, Measurable (f i X))
{b : ℝ} (hb : b 0) (hf : (i : ι) (z : Z), |f i z| b) :
μⁿ[fun ω : Fin n Ω i, |∑ k : Fin n, f i (X k)) - n μ[fun ω’ f i (X ω’)]|]
(2 * n) rademacherComplexity n f μ X := by
5.3 Normalized form: expected uniform deviation $\leq 2\mathfrak{R}_{n}$.
The file FoML/Main.lean packages a normalized statement for uniformDeviation, dividing the previous bound by $n$:
$\mathbb{E}_{\omega\sim\mu^{n}}\big[\mathrm{UD}_{n}(f;\mu,X;X\circ\omega)\big]\;\leq\;2\cdot\mathfrak{R}_{n}(f;\mu,X).$
This is the canonical “expected generalization gap bound” in Rademacher complexity theory.
FoML/Main.lean (excerpt)
theorem uniform_deviation_expectation_le_two_smul_rademacher_complexity
[Nonempty ι] [Countable ι] [IsProbabilityMeasure μ]
(hn : 0 < n) (X : Ω 甲)
(hf : i, Measurable (f i X))
{b : ℝ} (hb : 0 b) (hf : i x, |f i x| b) :
μⁿ[fun ω : Fin n Ω uniformDeviation n f μ X (X ω)] 2 rademacherComplexity n f μ X := by
5.4 Bounded-differences (Lipschitz) property of uniformDeviation.
To obtain high-probability bounds, the repo proves that changing one sample point changes uniformDeviation by at most $2b/n$ (the standard bounded differences constant) when all functions are bounded by $b$:
$\big|\mathrm{UD}_{n}(S)-\mathrm{UD}_{n}(S^{(i\leftarrow x^{\prime})})\big|\leq\frac{2b}{n}.$
FoML/BoundedDifference.lean (excerpt)
theorem uniformDeviation_bounded_difference
(hn : 0 < n) (X : Ω 甲) (hf : i, Measurable (f i X))
{b : ℝ} (hb : 0 b) (hf’: i z, |f i z| b)
(i : Fin n) (S : Fin n 甲) (x : 甲) :
|uniformDeviation n f μ X S - uniformDeviation n f μ X (Function.update S i x’)|
(n : ℝ)⁻¹ * 2 * b := by
5.5 McDiarmid tail bound and the “main theorem” (countable class).
Using a formal McDiarmid inequality (file FoML/McDiarmid.lean), FoML/Main.lean proves a tail bound around the mean and then combines it with the expectation bound to obtain the standard Rademacher-complexity tail estimate:
$\mathbb{P}\!\left(\mathrm{UD}_{n}\;\geq\;2\mathfrak{R}_{n}+\varepsilon\right)\;\leq\;\exp\!\left(-\frac{\varepsilon^{2}\,n}{2b^{2}}\right),$
where $|f_{i}(x)|\leq b$ uniformly. This is precisely the usual “symmetrization + McDiarmid” pathway found in standard textbooks and original papers on learning bounds.
FoML/Main.lean (excerpt)
theorem uniform_deviation_tail_bound_countable_of_pos
[MeasurableSpace 甲] [Nonempty 甲] [Nonempty ι] [Countable ι] [IsProbabilityMeasure μ]
(f : ι ℝ) (hf : i, Measurable (f i))
(X : Ω 甲) (hX : Measurable X)
{b : ℝ} (hb : 0 < b) (hf : i x, |f i x| b)
: ℝ} (hε : 0 ε) :
(μⁿ (fun ω 2 rademacherComplexity n f μ X + ε uniformDeviation n f μ X (X ω))).toReal
(- ε ^ 2 * n / (2 * b ^ 2)).exp := by
5.6 Extending main theorem from countable to separable index sets.
A recurring obstacle in mechanizing learning theory is measurability of suprema: mathlib provides Measurable.iSup for countable index types, but for uncountable parameter spaces measurability is not automatic. The development addresses this by separating proof obligations from modeling convenience. Theorem main_countable is proved for [Countable ι]; then main_separable lifts the result to the common case where $\iota$ is a separable topological space and the map $i\mapsto f_{i}(x)$ is continuous for each $x$.
The lifting works by reducing $\sup_{i\in\iota}$ to a supremum over a countable dense sequence. Formally, FoML/SeparableSpaceSup.lean proves a lemma (for real-valued continuous functions)
$\sup_{x\in X}g(x)\;=\;\sup_{n\in\mathbb{N}}g(\mathrm{denseSeq}(n)),$
stated as separableSpaceSup_eq_real. This lemma is then used to show definitional equivalence of the learning-theoretic quantities when replacing $f:\iota\to(\mathcal{X}\to\mathbb{R})$ by its restriction along denseSeq ι : Nat ι. In particular, FoML/Main.lean proves equalities
$\widehat{\mathfrak{R}}_{n}(\mathcal{F}\;x)=\widehat{\mathfrak{R}}_{n}(\mathcal{F}\circ\mathrm{denseSeq}\;x),\qquad\mathfrak{R}_{n}(\mathcal{F})=\mathfrak{R}_{n}(\mathcal{F}\circ\mathrm{denseSeq}),\qquad\Delta_{n}=\Delta_{n}^{\mathrm{denseSeq}},$
as empiricalRademacherComplexity_eq, RademacherComplexity_eq, and uniformDeviation_eq. The last equality is technically the most subtle: it requires continuity of the integral term $i\mapsto\int f_{i}(X(\omega)),d\mu(\omega)$, which is obtained via a dominated-continuity lemma under the uniform bound $|f_{i}(x)|\leq b$.
The final theorem main_separable applies main_countable to the countable family f denseSeq ι and rewrites the result back using the aforementioned equalities. Mathematically, it yields the same tail bound as main_countable but under topological assumptions ([SeparableSpace ι] plus continuity) rather than the purely set-theoretic [Countable ι] assumption.
FoML/Main.lean (excerpt)
theorem uniform_deviation_tail_bound_separable_of_pos
[MeasurableSpace 甲] [Nonempty 甲] [Nonempty ι]
[TopologicalSpace ι] [SeparableSpace ι] [FirstCountableTopology ι]
[IsProbabilityMeasure μ]
(f : ι ℝ) (hf : i, Measurable (f i))
(X : Ω 甲) (hX : Measurable X)
{b : ℝ} (hb : 0 < b) (hf : i x, |f i x| b)
(hf : x : 甲, Continuous fun i f i x)
: ℝ} (hε : 0 ε) :
(μⁿ (fun ω 2 rademacherComplexity n f μ X + ε uniformDeviation n f μ X (X ω))).toReal
(- ε ^ 2 * n / (2 * b ^ 2)).exp := by
6 Linear Predictors with $\ell_{2}$ and $\ell_{1}$ Regularization
To demonstrate the power of the abstract infrastructure, we specialize the Rademacher complexity to linear hypothesis classes with norm constraints.
6.1 $\ell_{2}$-bounded linear predictors ($O(1/\sqrt{n})$).
Let $d\in\mathbb{N}$ and consider the Euclidean space $\mathbb{R}^{d}$ (modeled as EuclideanSpace (Fin d)). For weights $w$ in an $\ell_{2}$ ball of radius $W$ and inputs $x$ in an $\ell_{2}$ ball of radius $X$, the class is
$f_{w}(x)=\langle w,x\rangle.$
The repo formalizes the standard empirical bound:
$\widehat{\mathfrak{R}}_{n}(\{f_{w}:\|w\|_{2}\leq W\};S)\;\leq\;\frac{XW}{\sqrt{n}}.$
The proof is recognizably the classical argument: rewrite the supremum via Cauchy–Schwarz, then compute the second moment of the Rademacher-weighted sum using sign orthogonality (all cross terms vanish). The orthogonality lemma is proved by an explicit sign-flip involution (rademacher_flip) in FoML/RademacherVariableProperty.lean.
FoML/Main.lean (excerpt)
theorem linear_predictor_l2_bound
[Nonempty ι] (d : ℕ) (W X : ℝ) (hx : 0 X) (hw : 0 W)
(Y : Fin n Metric.closedBall (0 : EuclideanSpace (Fin d)) X)
(w : ι Metric.closedBall (0 : EuclideanSpace (Fin d)) W) :
empiricalRademacherComplexity
n (fun (i : ι) a ⟪((Subtype.val w’) i), a⟫) (Subtype.val Y’)
X * W / √(n : ℝ) := by
6.2 $\ell_{1}$-bounded linear predictors ($O(\sqrt{\log(d)/n})$).
For $\ell_{1}$-bounded weights $\|w\|_{1}\leq W$ and $\ell_{\infty}$-bounded inputs $\|x\|_{\infty}\leq X_{\infty}$, standard learning theory gives
$\widehat{\mathfrak{R}}_{n}(\{x\mapsto\langle w,x\rangle:\|w\|_{1}\leq W\};S)\;\lesssim\;\frac{X_{\infty}W}{\sqrt{n}}\sqrt{2\log(2d)}.$
The repo proves essentially this statement (with explicit constants), using a Massart-type finite-class lemma for a $2d$-sized auxiliary class corresponding to signed coordinate projections; and a duality inequality reflecting $\sup_{\|w\|_{1}\leq W}\langle w,z\rangle=W\|z\|_{\infty}$ (implemented with careful absolute values and finiteness). Classical references typically attribute the finite-class step to Massart and standard concentration methods.
FoML/Main.lean (excerpt)
theorem linear_predictor_l1_bound
[Nonempty ι] (d : ℕ) (Xinf W : ℝ)
(hX : 0 Xinf) (hW : 0 W) (d_pos : 0 < d) (n_pos : 0 < n)
(Y : Fin n LinftyBall (d := d) Xinf)
(w : ι L1Ball (d := d) W) :
empiricalRademacherComplexity n
(fun i a => (∑ j : Fin d, (w i).1 j * a j))
(Subtype.val Y’)
(Xinf * W / Real.sqrt (n : ℝ)) * Real.sqrt (2 * Real.log (2 * d)) := by
7 Dudley Entropy Integral Bound
Beyond linear predictors, we formalized the Dudley entropy integral bound, which relates the Rademacher complexity of an arbitrary function class to its covering numbers under the empirical $L^{2}$ metric. This requires a substantial additional (half a year of) infrastructure layer, because Dudley’s theorem is naturally expressed in terms of metric entropy with respect to a sample-dependent pseudometric.
7.1 Empirical pseudo-metric and covering numbers.
For a fixed sample $S:\mathrm{Fin}\ n\to\mathcal{X}$ and functions $f,g:\mathcal{X}\to\mathbb{R}$, the repo defines an empirical $\ell_{2}$-type seminorm
$\|f-g\|_{S}:=\sqrt{\frac{1}{n}\sum_{k=1}^{n}(f(S_{k})-g(S_{k}))^{2}},$
implemented via empiricalNorm and empiricalDist in FoML/PseudoMetric.lean. This induces a pseudo-metric structure on the set of functions restricted to the sample.
Covering numbers are then defined for totally bounded sets in a pseudo-metric space:
$N(\varepsilon):=\min\{\,|T|:T\subseteq A,\ A\subseteq\cup_{t\in T}B(t,\varepsilon)\,\},$
implemented as coveringNumber ha ε using Nat.find on the finite approximation lemma for totally bounded sets (coveringNumber_exists) in FoML/CoveringNumber.lean.
7.2 The Dudley-type theorem as formalized.
The central statement in FoML/Main.lean is a wrapper around dudley_entropy_integral from FoML/DudleyEntropy.lean. It bounds the without-absolute-value empirical Rademacher complexity in terms of an entropy integral:
$\widehat{\mathfrak{R}}^{(\mathrm{no\ abs})}_{n}(F;S)\;\leq\;4\varepsilon\;+\;\frac{12}{\sqrt{n}}\int_{\varepsilon}^{c/2}\sqrt{\log N(u)}\,du,$
where $c$ is an upper bound on all empirical norms $\|F(f)\|_{S}$ and $\varepsilon<c/2$, and where $N(u)$ is the covering number of the class realized as a subset of EmpiricalFunctionSpace.
FoML/Main.lean (excerpt)
theorem dudley_entropy_integral_bound
{甲 : Type v} {n : ℕ} : Type u} [Nonempty ι] {F : ι ℝ} {S : Fin n 甲} {c ε : ℝ}
_pos : 0 < ε) (h : TotallyBounded (Set.univ : Set (EmpiricalFunctionSpace F S)))
(m_pos : 0 < n) (cs : f : ι, empiricalNorm S (F f) c)
_le_c_div_2 : ε < c/2) :
empiricalRademacherComplexity_without_abs n F S
(4 * ε + (12 / Real.sqrt n) *
(∫ (x : ℝ) in ε..(c/2), √(Real.log (coveringNumber h x)))) := by
7.3 Relation to classical Dudley bounds.
Dudley’s 1967 work [9] introduced entropy integrals to control sample regularity and suprema of Gaussian processes. Modern empirical-process theory adapts similar “chaining” ideas to Rademacher/Gaussian complexities, yielding inequalities of the form $\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\sum_{i}\sigma_{i}f(x_{i})\right]\lesssim\int\sqrt{\log N(\varepsilon)}\,d\varepsilon$. The Lean development follows this blueprint explicitly: it constructs dyadic radii $\varepsilon_{j}=c/2^{j}$, chooses finite $\varepsilon_{j}$-covers (via coveringFinset), defines a chaining approximation chainApprox, proves a telescoping decomposition, bounds the “tail” by $\varepsilon$ (Part A), bounds increments by a Massart finite-class lemma (Part B), and converts the resulting finite sum into an interval integral using a monotonicity-based Riemann-sum comparison.
8 Design Choices
We review the particularly challenging components of the development and the methodological choices that enabled progress.
8.1 Independent Variables
In the Rademacher complexity, the training dataset is assumed to be i.i.d.. This is essential, for example, in the symmetric arguments. In a textbook-style formulation, i.i.d. variables $X_{1},\ldots,X_{n}$ are formulated as a map $\mathrm{Fin}\,n\to\Omega\to\mathcal{X}$. However, we noticed that it is much convenient to formalize them as compositions of a single variable $X:\Omega\to\mathbb{R}$ with the coordinate projections $\Omega^{n}\to\Omega$, letting $\Omega^{n}$ as the base probability space. In this view, independence need not be explicitly assumed as it follows directly from the construction, which helps simplifying descriptions of theorems.
On the other hand, in the McDiarmid’s inequality, the random variables $X_{1},\ldots,X_{n}$ are assumed to be independent but need not be identically distributed. So we formalize them as distinct functions $X:\mathrm{Fin}\,n\to\Omega\to\mathbb{R}$ and explicitly assume independence as an explicit condition. Here the base probability space is $\Omega$.
The fact that the former construction $\Omega^{n}\to\Omega\to\mathcal{X}$ satisfies the independence condition required in the latter is a theorem that should be proved. We could not find this result in Mathlib, so we supplied our own proof. We also attempted another formalization that defines Rademacher complexity via probability mass functions, but deriving its properties along this path proved difficult, and we abandoned that approach.
8.2 Topological Details of Index Set $\iota$ and Hypotheses Class $\mathcal{F}$
The Rademacher complexity is an expectation of the supremum over hypotheses class $\mathcal{F}$, and in modern machine learning settings, $\mathcal{F}$ is often uncountable. When $\mathcal{F}$ is uncountable, however, measurability is not generally preserved under pointwise $\sup$. We therefore first prove results in the countable case and then extend to the separable case. Even when uncountable, a separable family allows the $\sup$ of continuous functions to be computed over a countable dense subset, reducing to the countable case. However, this requires that continuity be preserved under integration, which in turn requires first countability (a point that is, to our knowledge, not emphasized in standard textbooks). These conditions are satisfied in the spaces relevant to our applications, so they pose no practical obstacle.
8.3 Conditional Expectation
Textbook proofs of McDiarmid’s inequality typically proceed via conditional expectations. However, because conditional expectation is defined abstractly, it is difficult to carry out concrete computations directly from that definition. (At the time we were developing the formalization, Mathlib have not yet included the Doob–Dynkin lemma; now there is Mathlib.MeasureTheory.Function.FactorsThrough). We therefore avoided conditional expectations and instead defined the relevant quantities directly by integration, using independence. This changes the proof order: in the original argument the constructed sequence $Y$ is a martingale by construction, whereas in our approach (without conditional expectations) we establish the martingale property from independence.
8.4 Integral and Supremum
A large fraction of the lines in many lemmas is devoted to handling integrability conditions (and the treatment of $\sup$), which are often considered routine. Each time we performed an algebraic manipulation under the integral sign, we had to supply a fresh proof of integrability.
8.5 Dudley’s entropy integral without_abs.
The Dudley’s entropy integral bound is stated for empiricalRademacherComplexity_without_abs. In standard learning-theory presentations, one often places absolute values inside the supremum, or uses centered classes, to simplify symmetrization and contraction arguments. We instead provide a separate lemma showing without_abs is bounded by the absolute-value version under appropriate boundedness assumptions (file FoML/RademacherVariableProperty.lean).
8.6 Generative AI assistance
Our project spans more than two years of manual formalization. During the early stages, available large language models tended to produce Lean 3 code and were not practically useful. Since mid-2025 the models improved markedly and began producing syntactically correct Lean 4 code; we used generative AI to assist with routine yet tedious tasks such as rewriting expressions to fit Lean’s type system, managing casts between numeric types, and expanding algebraic identities. For example, writing the empirical Rademacher complexity as a measure-theoretic expectation required meticulous adjustments of types. While generative AI streamlined these technical steps, it could not reliably construct full proofs; we therefore designed the high-level strategy and wrote most proofs manually.
9 Literature Overview
Tables˜1 and 2 summarize close literature to this study.
9.1 PAC Learning and VC Dimension (Lean, Rocq)
Table 1: Formalization of Machine Learning Theory
Bentkamp et al. (2016,2019) [6, 7] Isabelle/HOL Expressive power superiority of deep over shallow
Bagnall and Stewart (2019) [4] Coq Generalization error bounds for finite hypothesis class
Tassarotti et al. (2021) [14] Lean PAC learnability of decision stumps
Vajjha et al. (2021) [17] Coq Convergence of reinforcement learning algorithms
Vajjha et al. (2022) [18] Coq Stochastic approximation theorem
Hirata (2025) [10] Isabelle/HOL No free lunch theorem
(ours) Lean Generalization error bound by Rademacher complexity
and Dudley’s entropy integral
Yuanhe et al. (2026) [21] Lean SLT with Gaussian Lipschitz concentration
and Dudley’s entropy integral
Table 2: Formalization of Concentration Inequalities
Markov/Chebyshev Lean (Mathlib), Rocq (MathComp-Analysis[1], IBM/FormalML[17]),
Isabelle/HOL (HOL-Probability)
Azuma-Hoeffding Lean (Mathlib)
McDiarmid (ours), Isabelle/HOL (AFP) [11]
Gaussian Lipschitz Lean [21]
A hypothesis class with finite VC dimension (or simply a finite class) admits Probably Approximately Correct (PAC) generalization bounds. Early formalizations focused on specific cases as follows:
Finite Hypothesis Classes:
Bagnall & Stewart (2019) [4] proved a general PAC bound in Rocq for any finite hypothesis class using Hoeffding’s inequality. Essentially, if $\mathcal{F}$ is finite, with probability $1-\delta$ the true error is within $\varepsilon$ of the training error for $n\gtrsim\frac{1}{\varepsilon^{2}}(\ln\|\mathcal{F}\|+\ln\frac{1}{\delta})$. Their Rocq development (part of the MLCERT system) used a union bound and Chernoff/Hoeffding bounds to link training and test errors. This formal result was applied to certify small neural network models’ performance. However, it was limited to finite $\mathcal{F}$ (not covering infinitely large model classes).
Decision Stump Class ($VC=1$):
Tassarotti et al. (2021) [14] gave a full Lean 3 proof that the concept class of decision stumps (threshold classifiers in $\mathbb{R}$) is PAC-learnable. This is a classic textbook example with VC dimension 1. The formal proof uncovered subtle measure-theoretic issues that are glossed over in informal proofs. For instance, textbooks often assume measurability of argmax operations on sample data without proof—the formalization had to rigorously prove measurability and proper probability space definitions for the learning algorithm. The authors structured the proof to separate combinatorial reasoning about the algorithm’s behavior from the analytic reasoning about probabilities. They employed the Giry monad (in Lean’s category theory library) to handle distributions, and ultimately derived the standard PAC guarantee for decision stumps.
9.2 Formalizing Machine Learning Theory
Statistical learning theory extends beyond classical generalization bounds.
Expressiveness of Deep Neural Networks:
Bentkamp et al. (2016,2019) [6, 7] formalized in Isabelle/HOL a theorem that deep networks can represent certain functions exponentially more efficiently than shallow ones. The formalization, simplified and generalized the original proof in 2016 within Isabelle’s logic. To support the proof, Bentkamp developed libraries for linear algebra (matrix ranks), multivariate polynomials, and even Lebesgue measure integration. The result is not about generalization error, but rather about the capacity/representation power of deep vs. shallow networks—nevertheless, it showcases the application of proof assistants to core theoretical ML questions. It also enriched Isabelle’s libraries with notions like tensor products and rank, which are useful in learning theory.
Convergence of Reinforcement Learning:
Vajjha et al. (2021) [17] verified the convergence of value iteration and policy iteration for Markov Decision Processes in Rocq. In a follow-up, they formalized a stochastic approximation theorem [18] useful for analyzing RL algorithms. These efforts, under the IBM FormalML project, build a bridge between learning theory and formal verification by proving probabilistic convergence properties in Rocq. They required heavy use of Rocq’s analysis libraries and bespoke techniques (e.g. coinduction for probabilistic processes).
Statistical Learning Theory Based on Empirical Process:
In early 2026, Yuanhe et al. [21] opensourced an AI-assisted formalization of statistical learning theory in Lean 4, focusing on empirical process theory. Starting from existing code such as lean-rademacher [^1] https://github.com/auto-res/lean-rademacher.git and CLT [^2] https://github.com/RemyDegenne/CLT.git, their system automatically generated a large number of definitions and proofs including Gaussian Lipschitz concentration and Dudley entropy integral for sub-Gaussian process.
9.3 Concentration Inequalities and Background Formalizations
Many learning-theoretic proofs rely on concentration of measure results and related probabilistic inequalities. Over the 2010s–2020s, these foundational results have been increasingly formalized, often as prerequisites for the theorems above:
Isabelle/HOL:
Its probability theory library (HOL-Probability) already included basic inequalities like Markov’s inequality, Chebyshev’s inequality, and exponential tail (Chernoff/Hoeffding) bounds by the late 2010s. In 2023, Karayel and Tan [11] contributed an AFP entry “Concentration Inequalities” which adds more advanced results. This includes Bennett’s and Bernstein’s inequalities (for sub-exponential random variables), Efron–Stein’s inequality (variance bound via variance decomposition), McDiarmid’s inequality (bounded differences), and the Paley–Zygmund inequality. Thanks to this, Isabelle/HOL now boasts one of the most extensive collections of concentration results, all formally proven. For example, the formal McDiarmid inequality in Isabelle was crucially used as a point of comparison for the Lean development.
Lean:
Lean’s Mathlib library [16] gained basic measure theory around 2019, including Lebesgue integration and statistical independence. By mid-2020s, Mathlib had formal proofs of Markov and Chebyshev inequalities. This study also introduced Hoeffding’s lemma (a result that a bounded zero-mean variable is sub-Gaussian) and then proved Hoeffding’s inequality as a corollary of Chernoff bound techniques. It also implemented McDiarmid’s inequality from scratch in Lean 4. In 2025, Degenne implemented Azuma-Hoeffding on Mathlib independently with this study. In 2026, Yuanhe et al. implemented Dudley’s entropy integral (with other concentration inequalities such as Gaussian Lipschitz) independently with this study.
Rocq:
Historically, Rocq’s standard library did not include measure-theoretic concentration results. But developments like MathComp-Analysis [1] have added some pieces. Using these, the IBM FormalML team [17, 18] have formalized basic inequalities such as Markov and Chebyshev in Rocq. Hoeffding’s and McDiarmid’s inequalities were not present in Rocq as of early 2020s except in special-case proofs (e.g. MLCert implicitly used a form of Hoeffding’s bound for finite samples). In 2025, Affeldt et al. [2] formalized concentration inequalities in the Rocq proof assistant, developing infrastructure for measure-theoretic reasoning and exponential bounds. Their work includes Hoeffding’s and Bernstein’s inequalities.
10 Conclusion
We presented a Lean 4 formalization of generalization analysis via Rademacher complexity, built on Mathlib’s measure-theoretic probability infrastructure. At the core is a mechanically checked pipeline that mirrors standard textbook proofs: (i) formal definitions of empirical and expected Rademacher complexity and of the uniform deviation functional, (ii) a symmetrization theorem that turns uniform deviations into sign-averaged suprema, and (iii) a bounded-differences argument combined with a fully formal McDiarmid inequality to obtain high-probability generalization bounds. A major methodological contribution is the systematic treatment of measurability obstacles: results are first established for countable hypothesis classes (where measurability of suprema is readily available), and then lifted to separable topological parameter spaces through a countable dense reduction, together with continuity-under-integration lemmas.
Beyond the abstract theorem, we demonstrated that this infrastructure supports nontrivial specializations. We formalized classical empirical Rademacher bounds for linear predictors under $\ell_{2}$ and $\ell_{1}$ norm constraints, and we mechanized a Dudley entropy integral bound using covering numbers and a chaining construction. These examples illustrate that once the general Rademacher pipeline is established in a proof assistant, many concrete bounds become “executable corollaries” obtained by instantiating the abstract objects and discharging model-specific analytic estimates.
Looking forward, we expect several extensions to be both practically useful and methodologically interesting: incorporating standard contraction inequalities for Lipschitz losses, connecting Rademacher complexity to Gaussian complexity and sub-Gaussian process tools, expanding libraries for covering-number estimates for common model classes, and improving automation for routine measurability and integrability obligations. More broadly, we hope this development contributes to a foundation in which modern learning-theoretic proofs can be stated and verified at a faithful, measure-theoretic level, enabling trustworthy reuse in formal verification and certified ML pipelines.
Acknowledgments
We would like to thank the mathlib community members, particularly Rémy Degenne and Yaël Dillies, for their valuable comments regarding formalization of Hoeffding’s lemma. We would like to thank Reynald Affeldt, Alessandro Bruni, Tetsuya Sato, Hiroshi Unno, and Yoshihiro Mizoguchi for having productive discussion on our formalization project. This work was supported by JSPS KAKENHI 24K21316, and JST Moonshot R&D Program JPMJMS2236, PRESTO JPMJPR2125, BOOST JPMJBY24E2, CREST JPMJCR25I5